### v1

# Latent Beam Diffusion Models for Decoding Image Sequences

Abstract
While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency in image sequences. Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent frames. We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding. Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions. To address beam search’s quadratic complexity, we integrate a cross-attention mechanism that efficiently scores search paths and enables pruning, prioritizing alignment with both textual prompts and visual context. Human evaluations confirm that our approach outperforms baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment. By bridging advances in search optimization and latent space refinement, this work sets a new standard for structured image sequence generation.
1 Introduction
Figure 1: BeamDiffusion models leverage search in the latent denoising space for the best sequence of denoised images, where $I_{n}$ represents the image at step $n$ and $S_{n}$ represents the text associated with step $n$.
Image diffusion models [42, 12, 43] have made significant advancements in generating high-quality images. These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [34, 29, 37]. While they excel at generating individual images based on specific text prompts, they face challenges when it comes to creating coherent sequences of images. In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across multiple steps. This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence. This dynamic can result in sequences that lack coherence and visual consistency over time [4].
We propose to utilize beam search as a potential solution to this problem. Beam search is a well-established technique in text generation tasks like machine translation [18, 47, 21, 36], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions. However, despite its success in other domains, it has been barely explored in our context, leaving its potential untapped. In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting text prompts or adding semantic cues.
In this work, we explore beam search for latent prior selection in image sequence generation. While previous research has used latent representations as priors to improve sequence consistency [4, 32], it remains unclear which specific representations yield the best overall sequence. We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences. As illustrated in Figure 1, our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
Due to the quadratic complexity of beam search algorithm, one needs to prune the search paths to keep a fixed number of candidate paths. To this end, we incorporate a cross-attention mechanism [23] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps. This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
We evaluated the effectiveness of our approach on three datasets: in the Recipes domain, and in the out-of-domain DIY Tasks and Visual Storytelling. Human evaluations show that our model outperformed baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
2 Related Work
Decoding methods are essential for enhancing the quality of large language models (LLMs). Techniques such as beam search, top-k and nucleus sampling are widely studied for their effectiveness in balancing computational cost with generating coherent, contextually appropriate outputs [40, 13, 9]. Beam search has also proven effective in tasks like machine translation [27], image captioning [45, 19, 1], and visual storytelling [15]. Building on these ideas, our work innovates by applying beam search to explore the latent space in diffusion models for image sequence generation.
Ensuring coherence remains a significant challenge in multimodal sequence generation [10, 24, 6, 20]. Some approaches, like AR-LDM [28], encode the context of caption-image pairs into multimodal representations to maintain consistency across outputs. Similarly, Rahman et al. [31] incorporates the history of U-net [35, 38] latent vectors to preserve narrative coherence. However, these methods often face challenges such as computational inefficiency and the risk of noise propagation over time. Other approaches, such as Gill [22], combine LLMs with image encoder-decoder systems. While effective for certain tasks, these models suffer from alignment issues, with retrieved or generated images failing to match the narrative context, and they too often require complex training. Our method avoids these pitfalls by direct conditional generation against the guiding text while explicitly maintaining coherence via the beam search.
Interpreting diffusion models has become a focal point in recent research. Pont-Tuset et al. [30] anchor each word in textual descriptions to specific regions of images. Tang et al. [44] analyze text-image interactions through cross-attention maps in the denoising module, offering insights into how textual prompts influence the generated images. Dewan et al. [7] enhance this interpretability by leveraging partial information decomposition to analyze how various components of a text prompt impact the generated image. In contrast, our beam-search approach focuses on generating images directly from textual input. Rather than mapping text to pre-existing images, it generates image sequences that are visually coherent and are consistently aligned with the evolving sequence of textual narrative.
Bordalo et al. [4] propose to use latent information from previous steps to guide coherent generation of subsequent images. However, their method relied on a greedy heuristic-based selection. CoSeD [32] improved on this idea and introduced a contrastive learning approach [5] to rank images generated from different latent seeds. Yet, both approaches focused on local refinements rather than optimizing across the entire sequence. Our beam search approach lifts these local selection criteria to perform global optimization.
3 Synthesis of Non-Linear Visual Sequences
In the following section, we define the challenges in image sequence generation, recall the core principles of LDM, Sec. 3.1, and then we detail BeamDiffusion key aspects in terms of diffusion beams (Sec. 3.2.1) and pruning (Sec. 3.2.2) with a contrastive approach (Sec. 3.2.3).
3.1 Problem Formulation
Generating a full sequence of visually coherent images from textual scene descriptions presents a challenge. Given a sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each scene $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps.
Latent Diffusion Models (LDMs) [34] generate high-quality images in a compressed latent space. Instead of directly operating in pixel space, LDMs encode an input image $I$ into a latent representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [2]. The U-Net backbone [35], denoted as $\epsilon_{\theta}$, incorporates a cross-attention mechanism [17] to condition the diffusion process on an external input $\tau_{\theta}(s)$, where $s$ represents a textual prompt. LDMs are trained by minimizing:
$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right].$ (1)
While effective, LDMs generate images independently, starting from a randomly initialized latent $z_{T}$. This lack of temporal coherence makes it challenging to ensure consistency across generated sequences, especially in non-linear narratives [41], where maintaining character identity, lighting, and environmental details are crucial.
(a) The beam denoising search tree.
(b) Example of one complete decoded denoising process using contextual captions.
Figure 2: The BeamDiffusion model works in the denoising latent space by denoising the initial seed to latents and them decoding different beam paths. This shows how our method evolves in the latent space and how it can explore the latent space.
3.2 Beam Diffusion Model
To address the limitations of standard LDMs in generating coherent image sequences, we introduce the BeamDiffusion model. The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space. BeamDiffusion incorporates latent representations $\mathcal{Z}_{j}$ from previously generated images. These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence. This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the new scene description. In the following sections, we will detail the proposed method.
3.2.1 Latent Space Exploration with Diffusion Beams
To maintain both diversity and coherence in the generated image sequence, we propose a diffusion-based exploration of the latent space across multiple steps. At each generation step $j$, the set $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots,I_{j}^{(B)}\}$ consists of all possible candidate images, with $I_{j}^{(b)}$ representing the image generated at step $j$ for beam $b$.
Sampling the Denoising Latent Space.
In the first step of the sequence, $s_{1}$, the model generates multiple candidate images using different random seeds. This approach allows for a broader exploration of the latent space and reduces Bayes risk [3], as described in Appendix B. Hence, in this initial step, $s_{1}$, we obtain the first exploration samples from the denoising latent space, denoted as $\mathcal{Z}_{1}=\{z_{1}^{(1)},\ldots,z_{1}^{(K)}\}$.
Building on prior work [4, 32], incorporating latent denoising from earlier steps ($s_{<j}$) can enhance the denoising process. We define $\mathcal{Z}_{j}$ to be the accumulation of latent representations from all preceding images up to step $j$:
$\mathcal{Z}_{j}=\bigcup_{i\in\{j-n,\ldots,j-1\}}\{z_{i}^{(1)},\dots,z_{i}^{(K)}\},$ (2)
where each latent representation $z_{i}^{(k)}$ corresponds to the $k$-th iteration of the diffusion process for image $I_{i}$. By utilizing latents from $n$ previous stages (steps back), our approach ensures contextual consistency across generations.
Considering one particular search beam $b$ in the latent space, to generate a new candidate image $I_{j}^{(b)}$, we condition the diffusion process on a latent representation $z$ selected from the accumulated latents in $\mathcal{Z}_{j}$, as represented in Figure 2(a). This conditional dependency is captured by the following probability expression:
$\small p\left(I_{j}^{(b)}\mid\theta,z\in\mathcal{Z}_{j},s_{j}\right),$ (3)
where $\theta$ are the model parameters. This formulation ensures that all denoising processes are correlated, as each new image $I_{j}^{(b)}$ is generated based on latents $z$ drawn from $\mathcal{Z}_{j}$. In this way, each iteration in the denoising process acts as a seed for the next, establishing a dependency between successive denoising processes (Figure 2(b)).
Diffusion Beams.
Following the idea of beam search, which maintains a set of the most promising candidates at each step and extends them sequentially [46], we construct diffusion beams by iteratively expanding candidate images. Formally, for each step $j$, we generate the set of all candidate images $\mathcal{I}_{j}$ to compute the set of candidate diffusion beams $\hat{B}_{j}$. At the initial step $j=1$, there are no previous beams to combine, so the initial set of beams can be represented as,
$\small\hat{B}_{1}=\bigcup_{k}\{I_{1}^{(k)}\mid I_{1}^{(k)}\in\mathcal{I}_{1}\},$ (4)
where $\mathcal{I}_{1}$ is the set of initial candidate images generated using different random seeds indexed by $k$. For subsequent steps $j>1$, each beam $\hat{B}_{j-1}^{(b)}$ from the previous step is extended with new candidate images $I_{j}^{(b\cdot k)}$ from the set $\mathcal{I}_{j}$, and this extension is represented by the operator $\oplus$. The new set of beams is now represented as,
$\small\hat{B}_{j}=\bigcup_{(b,k)}\left\{\hat{B}_{j-1}^{(b)}\oplus I_{j}^{(b\cdot k)}\mid I_{j}^{(b\cdot k)}\in\mathcal{I}_{j}\right\}$ (5)
where $b$ is the index of a beam from the set $\hat{B}_{j-1}$, and $k$ is the $k$-th latent variable from $\mathcal{Z}_{j}$. For each combination of a previous beam $\hat{B}_{j-1}^{(b)}$ and latent $\mathcal{Z}_{j}^{(k)}$, the model considers a new candidate image $I_{j}^{(b\cdot k)}$ from the set of candidate images $\mathcal{I}_{j}$. The union over $(b,k)$ combines all possible sequences of previous beams with the new candidate images, expanding the search space for the next step of generation.
3.2.2 Pruning Diffusion Beam Paths
Tracking diffusion beams helps maintain context and coherence across the growing image sequence by exploring multiple potential continuations. However, as the sequence progresses, focusing on the most promising paths within the beam budget (the number of beams that can be maintained simultaneously) becomes increasingly important to ensure the coherence of the generated images and optimize computational resources. To achieve this, our model prunes the diffusion beam paths, narrowing down the search space for the best candidates while promoting diversity and consistency in the maintained beams.
To prune candidate beams, the model scores each candidate $\hat{B}_{j}$ by the likelihood of image $I_{j}$ in the beam given the previous images $I_{<j}$ and the step text. This is represented as $p(I_{j}|s_{j},B_{j-1}^{(b)})$, where $B_{j-1}$ represents the set of the most promising candidate beams from the previous step. Formally, the scoring is defined as:
$\small score(\hat{B}_{j}^{(b)})=\sum_{i=1}^{j}\log p(I_{i}|I_{1},\dots,I_{i-1},s_{1},\ldots,s_{i}),$ (6)
where we sum the individual log-probabilities of each image in the sequence, reflecting the overall coherence and alignment with the context set by the previous images and prompts.
After scoring all candidate images in $\mathcal{I}_{j}$, the search space is refined by performing a beam search operation. This operation selects the most promising $w$ diffusion beams, effectively limiting the beam width to $w$. The refinement process is expressed as,
$\small B_{j}=\mathop{\mathrm{argmax}}_{B_{j}\subseteq\hat{B}_{j},|B_{j}|=w}\left(\sum_{i=1}^{w}score(\hat{B}_{j}^{(i)})\right)$ (7)
where, $B_{j}$ represents the most promising $w$ candidate diffusion beams at the step $j$, ensuring that the updated beam evolves to explore the most promising options, guided by the previous context and the newly generated candidates, while maintaining a controlled beam width.
To maintain a broad search space in the early stages, pruning is delayed until after the second step. During this initial phase, all candidates are retained for exploration. After the second step, pruning is applied, limiting the candidates to the $w$ most promising. This approach balances exploration and computational efficiency, focusing the model’s efforts on the most promising candidates in the later stages of generation.
Figure 3: Visualization of cross-attention in latent diffusion, showing how BeamDiffusion focuses on different regions of the image based on the prompt at each stage. The heatmap highlights the areas with the strongest attention.
3.2.3 Contrastive Score
While the basic pruning mechanism outlined above provides a foundation for selecting candidate images based on raw probabilities, the scoring function is now implemented as contrastive classifier. This classifier improves the image selection by learning from past sequences and aligning the candidates with the evolving context [32]. As illustrated in Figure 3, the heat maps show how BeamDiffusion attends to different regions of the input text at various stages, highlighting the evolving alignment between textual prompts and image features.
To compute these probabilities, we refine the approach using a softmax-based method to determine the likelihood of each candidate image $I_{j}$ relative to all other possible candidates in the sequence. In this refined process, the probability $p(I_{j}|s_{j},B_{j-1}^{(b)})$ is replaced by the output of the classifier $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$, which approximates the likelihood of each candidate fitting the sequence. The more detailed version of the scoring mechanism can be expressed as,
$\scriptsize\begin{split}score(\hat{B}_{j}^{(b)})&=\sum_{i=1}^{j}\log\left(\frac{\exp\left(\varphi(I_{i},s_{i},B_{i-1}^{(b)})\right)}{\sum_{I_{i}^{\prime}\in\mathcal{I}_{i}}\exp\left(\varphi(I_{i}^{\prime},s_{i},B_{i-1}^{(b)})\right)}\right),\end{split}$ (8)
where the denominator in the softmax function normalizes the scores by summing over the candidate images $I_{i}^{\prime}$ for beam $b$ within the set $\mathcal{I}_{i}$. This ensures that the score of selecting a particular candidate image $I_{j}$ is computed relative to all other candidate images in the sequence, capturing how well it fits the overall sequence in terms of visual content and textual alignment.
Training.
To train the classifier effectively, we employ a contextual sequence training approach, which enables the model to capture sequential relationships across different tasks. During training, the true label $l_{t,j}$ indicates whether the candidate image $I_{j}$ fits correctly into the sequence, given the previous steps $B_{j-1}^{(b)}$ of the $b$-th beam. Here, $t$ refers to the task index and $j$ represents the step index within the sequence. The model’s prediction $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$ is based on the candidate image $I_{j}$, its textual description $s_{j}$, and the context provided by the prior steps. The model aims to minimize cross-entropy loss by aligning the candidate image with its context, ensuring it fits well within the sequence according to the true labels $l_{t,j}$, defined as:
$\small\operatorname*{argmin}\sum_{t=1}^{M}\sum_{j=1}^{N}l_{t,j}\log(\varphi(I_{j},s_{j},B_{j-1}^{(b)}))$ (9)
Through contrastive training, the model learns to align representations of similar steps while enhancing the distinction between steps that differ semantically or visually.
3.2.4 Contextual Scene Descriptions
In many cases, step descriptions alone may lack context as precise visual prompts. These step descriptions are often interdependent, relying on prior steps for contextual understanding, or may miss visual details, e.g. ”Put the ice cream in freezer for 1 hour”, or actions that involve multiple tasks, such as ”Slice the tomatoes, grate the cheese, and stir the sauce”. To address this challenge, we leverage Gemini [33] to refine step descriptions into visually detailed prompts while considering sequence context. Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (10)
Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps. For further information, see Appendix E.
Recipes DIY VIST
Method Human (%) Gemini (%) Human (%) Gemini (%) Human (%) Gemini (%)
Greedy/CoSeDlen=1 [32] 13.3 16.7 20.0 10.0 27.2 16.7
Nucleus Sampling [14] 33.3 16.7 30.0 30.0 18.2 26.6
GILL [22] 0.0 3.3 10.0 20.0 0.00 0.0
BeamDiffusion 53.3 63.3 40.0 40.0 54.5 56.7
Table 1: Frequency of method selection by human annotators and Gemini.
4 Experimental Setup
In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences. We outline the datasets used, identify other baseline methods, and explain the human and Gemini evaluations focused on semantic and visual consistency across generated sequences. All data and implementations will be released after publication.
Datasets.
We used a dataset consisting of publicly available manual tasks in the Recipes and DIY domains [4], as well as the Visual Storytelling (VIST) dataset [16]. Recipes serve as the in-domain dataset, while DIY tasks and VIST are out-of-domain. These out-of-domain datasets allow us to assess the generalizability of our model across different task domains. More details can be found in Appendix D.
Baselines.
To assess the performance of our image sequence generation method for multistep tasks, we compared BeamDiffusion to two other decoding approaches, as well as GILL [22], focusing on different strategies for generating consistent image sequences:
  • Greedy/CoSeDlen=1 [32]: In this method, a new image is generated at each step using the first five latents from the previous image, selecting the best overall image. Although similar to CoSeDlen=1, it relies solely on CLIP without alterations, whereas CoSeD uses a trained module for selection.
  • Nucleus Sampling [14]: Rather than choosing the single most likely image, nucleus sampling randomly selects from the set of images whose cumulative probability exceeds a specified threshold $p$, according to CLIP. This allows for more diversity in the generated sequence while still emphasizing high-probability outcomes.
  • GILL [22]: Unlike the other methods, GILL does not explicitly search the latent space. Instead, it ensures coherent image sequences by iteratively generating images, while maintaining consistency across the sequence. GILL achieves this by deciding when to retrieve an image from a candidate set or when to generate a new image, based on the context of the previous images and text inputs.
Human Evaluation.
In the first task, the annotators compared pairs of beam search configurations in a ”battle-style” format to select which beam configuration was the best regarding beam width ($w$ in 7) and steps back ($n$ in 7). The winning configuration is compared against other configurations until a new configuration is selected as better. The selection criteria focused on semantic alignment and visual consistency. Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
Having chosen the optimal configuration, we need to assess how it compares to other methods. In the second annotation task, BeamDiffusion was compared with our baseline models. The annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency, with the possibility of choosing multiple options or none. For further information, see Appendix C.
Gemini Evaluation.
To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines. Its selection process closely mirrored that of human annotators, with one key difference: Gemini tends to default to ”both good” when given the option. To avoid this, we removed the ”both good” and ”both bad” options from its evaluation to ensure more accurate results.
We also refined our method for determining the best generation technique. Initially, presenting all four sequences at once caused the model to focus primarily on the last two, leading to biased results. To address this, we switched to pairwise evaluations, comparing sequences two at a time, which provided a more balanced and reliable assessment. For details, please refer to Appendix C.
Human-Gemini Agreement.
We conducted an alignment assessment between human annotators and Gemini. We asked 10 annotators to rate 16 Recipes, 10 DIYs, and 10 VIST tasks. After this assessment, we computed Fleiss’ kappa [8], obtaining a value of 70%, indicating a substantial agreement. Given this substantial alignment, we extended the evaluation to the remaining sequences with Gemini.
Steps Back Beam Width
2 3 4 6
2 8.33 8.33 25.00 8.33
3 16.67 - - -
4 16.67 8.33 8.33 -
Table 2: Influence of different hyperparameter values according to human annotations.
Steps Back Beam Width
2 3 4 6
2 5.71 2.86 48.57 5.71
3 28.57 - 2.86 -
4 - - 2.86 -
Table 3: Influence of different hyperparameter values according to Gemini annotations.
Figure 4: Non-linear steps selection, according to selection opportunities.
5 Results and Discussion
In the following sections, we demonstrate the performance of BeamDiffusion and examine different aspects: hyperparameters, non-linear decoding properties and a qualitative analysis.
5.1 General Results
The results in Table 1 compare the performance of four methods, Greedy/CoSeDlen=1, Nucleus Sampling, GILL, and BeamDiffusion, evaluated through human annotations and the Gemini model across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain).
Across domains, BeamDiffusion consistently outperforms other methods. In the Recipes domain, it leads with 53.3% in human evaluation and 63.3% in Gemini. Nucleus follows with 33.3% (human) and 16.7% (Gemini), while greedy scores 13.3% (human) and 16.7% (Gemini). In DIY, BeamDiffusion achieves 40.0% in both evaluations, again outperforming greedy and nucleus. In VIST, BeamDiffusion leads with 54.5% (human) and 56.7% (Gemini), with greedy and nucleus scoring significantly lower. GILL consistently underperforms, receiving 0% selection in VIST across both evaluations. In Recipes, it scores just 3.3% in Gemini and remains the weakest performer across all methods. While it performs slightly better in DIY, it still lags behind other approaches. Although BeamDiffusion excels in Recipes and VIST, performance in DIY remains more challenging, with smaller performance gaps between methods. This reflects the inherent complexity of the DIY domain, where generating coherent outputs is particularly difficult.
Overall, BeamDiffusion stands out as the most effective method across all domains, particularly in Recipes and VIST. However, its performance in DIY highlights the need for a more effective search in the latent space to ensure consistency and coherence across diverse tasks.
Figure 5: Example of the Recipes domain task ”Jamaican Callaloo With Shrimp”, with a sequence of 4 steps.
5.2 Best Beam Search Configuration
The hyperparameters beam width ($w$ in Eq. 7) and steps back ($n$ in Eq.  2), play a crucial role in the performance of BeamDiffusion. In this section, we evaluated different configurations across four beam widths (2, 3, 4, and 6) and three steps back (2, 3, and 4). In this study, we specifically focus on latents indexes 0, 1, 2, and 3, denoted as $K$ in Eq. 2, which yield the most promising results. A more detailed discussion on the impact of latents is provided in Appendix A.1.
We first performed human annotations on six tasks to identify the optimal beam search configuration, see Appendix C for details. The results in Table 3 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%. Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed competitive performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration. The configurations with beam width 6 performed the worst in all the settings tested.
To assess the reliability of these findings, we used Gemini for the same evaluation on six tasks, achieving an 85% Fleiss’ Kappa agreement [8], indicating near-perfect alignment. The results from the extended Gemini evaluation (Table 3) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained the best, now achieving a 48.57% selection rate, a significant improvement over the 33.33% from the initial six-task evaluation. Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative. However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
5.3 Non-Linear Sequence Denoising
To analyze how beam diffusion explore the latent space as formulated in Sec. 3.2.1, we examined the step from which the latent representation was selected during the generation process. For an analysis of latent selection, see Appendix A.2. We focused the step selection process in four-step sequences (Figure 4). Step 1 was chosen 69.9% of the time, highlighting its key role in shaping the sequence. Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%. The sum of these percentages exceeds 100% because they represent the relative selection frequency of each step, considering the number of possible selection opportunities. Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate. Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total. This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate stages. By balancing local accuracy with global consistency, BeamDiffusion ensures high-quality image sequences that remain visually and contextually aligned.
Figure 6: Example of the out-of-domain DIY task ”How to Make a Door”, with a sequence of 5 steps.
Figure 7: Example of the out-of-domain VIST story ”July 4th Barbecue”, with a sequence of 5 steps.
5.4 Qualitative Analysis
To illustrate how different decoding methods impact the semantic and visual consistency of generated output, we present several examples in Figures 57, and in Appendix F.2. In Figure 5, we see that BeamDiffusion consistently maintains the same pan across all steps, a quality not always achieved by greedy or nucleus methods. Meanwhile, GILL struggles to follow the text from step to step, leading to more inconsistencies in the outputs.
Looking at the out-of-domain DIY and VIST datasets, in Figure 6, we observe a task with intricate details. Although other methods fail to consistently capture the color of the plywood, BeamDiffusion outperforms by maintaining this detail, demonstrating its strength in preserving fine-grained elements in complex tasks. In Figure 7, we again observe significant inconsistencies with the greedy and nucleus methods, which alternate between black-and-white and color images. In contrast, BeamDiffusion keeps the color consistent throughout the sequence. Additionally, background elements, such as the bushes, remain stable across the first three steps. Even the persons in first step closely resemble those in the third, further emphasizing the BeamDiffusion consistency. At the same time, GILL consistently struggles to follow the text, failing to capture the necessary details.
This analysis highlights BeamDiffusion ability to maintain high levels of semantic and visual consistency, even in out-of-domain tasks, much like its performance on Recipes.
6 Conclusions
In this paper, we proposed a novel method to explore the latent diffusion space with a beam search strategy, to generate image sequences. BeamDiffusion is a theoretically sound framework, drawing on the strengths of text-based decoding techniques [40], with clear experimental gains over alternative methods, as we demonstrated in generating image sequences with enhanced coherence and consistency.
Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences. The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity. In contrast, GILL, the only method in our study that does not perform latent search, struggles to maintain consistent alignment with the text instructions, resulting in more inconsistent and less coherent image sequences. Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process.
Appendix A Impact of Latent Selection on Beam Search Performance
Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search. By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence. To investigate this, we evaluate 15 tasks, each consisting of four steps, allowing us to systematically assess the effects of latent selection across a diverse set of sequences. Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
A.1 Impact of Latent Groups on Performance
Annotations.
Following the initial analysis of beam search configurations (Sec. 5.2), we further investigated the impact of using different latent configurations on the generated sequences. While latents 0, 1, 2 and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance. For these latents, the beam search configurations yield very limited selection rates, indicating that they are rarely selected in both human and Gemini annotations. Specifically, for human annotations evaluated across 6 tasks, the only meaningful result is from the configuration with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%. Other configurations show no selection rates.
Figure 8: Impact of using 1, 3, 5, and 7 latent configurations on image generation.
Similarly, for Gemini annotations evaluated on the same 6 tasks, the configurations do not yield any notable results, with no selection rates. When the evaluation is extended to 35 tasks, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%. Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [32, 4].
Automatic Metrics.
Given the negative impact of higher latents on performance, as observed in human annotations, we decided to investigate the underlying reasons for this trend. Specifically, we explored how latents affect text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [44]. DAAM helps interpret text-to-image diffusion models by analyzing the cross-attention maps within the denoising process. It generates heat maps that highlight how much each token in the prompt influences different parts of the generated image, offering a visual representation of text-image alignment. By applying DAAM, we aimed to determine whether higher latents distort or weaken the contribution of key tokens during generation.
We then used DAAM to analyze the cross-attention between the input tokens and the generated image. DAAM enables us to track how much each remaining token influences specific regions of the image by calculating the attention patterns in the model. The resulting heat maps visualize this influence, with the intensity of the heat maps indicating the level of contribution each token has on different parts of the generated image. This approach helps us understand how well the model attends to key parts of the prompt and ensures that the most meaningful tokens are appropriately reflected in the image.
To quantify text alignment, we calculated the average heat map maximum intensity for all remaining tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image. Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence. While this serves as an objective metric for evaluating text adherence, it remains a partial measure, as it focuses solely on text alignment without accounting for other important aspects of image quality, such as coherence, realism, or artistic fidelity. Figure 9 shows the average impact of different latent groups on text alignment. The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges. Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523) and the 8–11 range (1.435). In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores. This suggests that earlier-stage latents play a crucial role in preserving text fidelity, whereas later-stage latents may dilute or distort the contribution of key tokens.
Figure 9: The average influence of latent indices on text alignment across task data.
In Figure 15, we analyze text alignment across multiple steps using heat maps corresponding to different latent groups, revealing how the connection between the input text and the generated image evolves throughout the process. Certain areas, such as the background and the bowl, show little to no heat map activation, indicating that these regions are less influenced by the text and more shaped by the model’s learned priors and overall scene composition. Since the model generates new images by building on latents from previous steps, these early decisions about text relevance persist throughout the process, influencing later refinements. Additionally, when examining generations using later latent groups, we find that this information remains more stable across all steps. This suggests that latents at these stages carry more structural and stylistic details, focusing on consistency rather than continuous adjustments based on the input text.
While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts. For example, later latent groups (4-7) maintain stability in visual elements throughout the process, preserving previously established features. However, this comes at the expense of adapting to new or changing textual instructions. In contrast, early latent groups (0-3) demonstrate strong alignment between the text and the generated image, with consistently high heat map intensity across all steps. This highlights their role in grounding the image in the input text. For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
A.2 Latent Selection on Image Sequence Generation
Figure 10: The average number of times that the model selected a given latent to generate the next image.
Following our evaluation of the impact of latent groups on performance, we now examine the influence of individual latents within the groups that exhibit better text alignment (0-3). As shown in Figure 10, Latent 3 is used most frequently, accounting for 44.4% of selections, followed by Latent 0 at 20%. In contrast, Latents 1 and 2 are used less often, with Latent 1 at 11.1% and Latent 2 at 8.9%. This uneven distribution suggests that Latents 0 and 3 are particularly effective at capturing key features at critical points in the generation process, which beam search exploits to enhance sequence coherence. The ”Rand seed,” which introduces variability into the process, accounts for 15.6% of selections. This highlights beam search’s flexibility, allowing it to explore a range of possible outputs while preserving the overall structure of the image sequence.
In general, these findings reinforce the importance of latent selection strategies in optimizing image generation, suggesting that models could benefit from prioritizing the most impactful latents to achieve both coherent visual sequences and stronger text-image alignment.
Figure 11: Heat maps illustrating cross-attention in latent diffusion across different decoding methods. The visualizations show variations in how each approach preserves or redistributes attention to textual cues throughout the generation process.
Figure 11 illustrates the impact of different latent search methods on text alignment. The heat maps reveal how attention patterns shift based on the chosen decoding strategy, highlighting the differences in how each approach preserves or redistributes focus on textual cues throughout the image generation process. As shown, particularly in step 3, BeamDiffusion demonstrates superior text alignment compared to the other methods. Not only does it maintain a stronger connection to the input text, it also ensures better overall sequence coherence. This improved alignment and consistency across generations are key advantages of BeamDiffusion.
These variations provide further insight into how latent search methods can influence both the coherence of the generated sequence and the alignment between the input text and visual output.
Appendix B Minimizing Decoding Risk
The generation process using diffusion models begins with a seed that determines the initial conditions of the model. The choice of this seed significantly affects the outcome, as it influences the path the model takes during generation. Using multiple random seeds in the first step ensures a diverse exploration of possible outputs, avoiding premature narrowing of options into a single path based on that seed. This limits the diversity of the results and increases the chance of missing better options.
Greedy Decoding and Nucleus Sampling.
In both greedy decoding and nucleus sampling, we use the same approach in the first step. Multiple random seeds are employed to generate a set of candidate images. Each candidate is then evaluated using the CLIP score mechanism, which compares the generated images with the step text. The image that best aligns with the text, according to the CLIP score, is selected to continue the process. This method ensures that both greedy decoding and nucleus sampling benefit from multiple diverse starting points, increasing the quality and variety of the generated outputs.
BeamDiffusion.
In BeamDiffusion, relying on a single random seed in the first step effectively makes the process a greedy selection, as it locks the generation into a specific path. This limits diversity and reduces the chances of discovering higher-quality outputs. By using multiple random seeds, each seed generates a different decoding trajectory, helping to broaden the search space and increasing the likelihood of finding better solutions. Using multiple random seeds in the initial step enhances the exploration of the latent space, ensuring better and more diverse results across different decoding strategies.
Random Seeds for mid-sequence Generation.
The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{l}\}$ may have one scene in the middle of the sequence that should be independent. To address this, for steps $j>1$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
Appendix C Annotation Tasks
Human annotations.
To facilitate the selection of the best beam search configuration, we developed a custom website specifically for human annotation. We found that existing tools lacked the flexibility necessary for our specific use case, especially when evaluating multiple configurations simultaneously. To select the best beam search configuration, 5 annotators are presented with 24 different sequences, each corresponding to a unique beam search configuration.
In total, the 24 configurations were generated by combining three key factors:
  • Steps Back: This refers to how many previous steps the model uses from its latent representations to explore new candidate outputs. We evaluated three different settings for the number of steps back: 2, 3, and 4, which correspond to using latents from the last two, three, or four images, respectively.
  • Number of Latents: We tested two sets of latent configurations {1, 3, 5, 7} and {0, 1, 2, 3} to explore how varying levels of latent detail impact the quality of the generated sequences.
  • Beam Width: We explored four different beam widths {2, 3, 4, 6} which determine how many candidate sequences are retained at each step during the search process.
Evaluating all of these at once proved to be exhausting and difficult to manage. To simplify the evaluation process, we adopted a ”battle-style” annotation method. In this approach, 5 annotators compare two configurations at a time and select the one they find has better overall coherence. If neither configuration stands out, annotators can select ”both good” or ”both bad” (Figure 12), in which case the computationally cheaper option advances to the next round. This process is repeated until we identify the best configuration. This method focuses on direct comparisons, making it easier to pinpoint the most effective beam search settings while also considering computational efficiency.
Figure 12: Example of the human annotation battle round for selecting the best beam search configuration.
To evaluate all models, 10 annotators are presented with four generated sequences side by side: one from the top-performing beam search configuration (selected during the initial annotations) and three others from alternative decoding methods, such as greedy search, nucleus sampling, and GILL. Figure 13 illustrates an example of this comparison process, where the annotators assess the sequences to determine the most effective approach.
Figure 13: Example of a human annotation to choose the best decoding method.
Gemini evaluation.
Figure 14 shows the prompt used in Gemini to annotate the best configuration and the best method. This prompt was used to guide the evaluation of image sequences based on visual and semantic consistency.
Gemini Prompt: ”Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames. Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements. Semantic Consistency: Images must logically follow the described steps, clearly and accurately. Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic). Pay close attention to even minimal differences, as the sequences are very similar. If both sequences are equally good or bad, choose the one that is slightly better. Choose exactly one of the following options: - ”First sequence” - ”Second sequence” ”
Figure 14: Gemini prompt used to select the best beam search configuration and best method.
Appendix D Datasets
Recipes and DIYs.
Each manual task in both the Recipes and DIYs datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps. Most tasks feature an image for each step. The DIYs dataset includes a variety of tasks, from simpler projects to more complex ones, with detailed steps.
VIST.
We used the Descriptions of Images-in-Isolation (DII) annotations, as they allow for the generation of more precise and informative visual prompts. These annotations provide standalone descriptions of images, making it easier to capture their key visual elements without relying on surrounding context. The dataset includes 50 000 stories, offering a diverse range of visual and textual information. To ensure the most accurate and relevant annotation for each image, we leveraged CLIP. By using CLIP, we were able to select the annotation that best represented the image, improving the quality of our generated prompts and enhancing the overall alignment between text and visuals.
Appendix E Challenges
In the development and evaluation of the BeamDiffusion method, we encountered several challenges that required careful consideration and innovation. These challenges spanned multiple aspects of the model, ranging from the difficulty of generating accurate and concise captions, to identifying suitable domains for testing, and addressing the computational complexity of the generation process. While we were able to propose solutions for some of these challenges, others remain open and present ongoing areas of research. In this section, we outline the key challenges we faced and the strategies we employed, while acknowledging that not all of them have been fully addressed.
Captions.
One of the challenges encountered in this process is generating captions that accurately describe the steps involved in the sequence. Often, a single step may consist of multiple sub-steps, each contributing to the overall generation. This makes it difficult to write a single concise caption that captures the full scope of each step. The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each stage but also conveys the progression and complexity of the entire sequence. Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
Domains.
Identifying a domain that effectively highlights the advantages of the BeamDiffusion method, particularly the benefits of beam search, is still a challenge. While BeamDiffusion provides improvements in consistency and coherence, not all domains clearly showcase these advantages. Certain domains may lack the complexity or diversity needed to demonstrate the method’s strengths, such as its adaptive latent selection or its ability to maintain alignment with contextual instructions. Finding a domain that not only aligns with the method’s capabilities but also provides a clear and compelling demonstration of its improvements remains an ongoing challenge.
Complexity of the generation.
The Beam Diffusion model generates multiple candidate images at each step by exploring a range of latents. We analyze the complexity of the model for a steps back of $n$ steps and $k$ latents explored. In the initial steps, the complexity is high because pruning begins only after the second step. For each beam, $k\times n$ latents are explored, and the search space grows exponentially, resulting in significant computational cost in the early stages. During the first two steps, the complexity grows with the number of candidates and latents, proportional to $O((k\times n)^{j})$ for $j\leq 2$. Specifically, for $j=1$, the complexity is proportional to the number of explored random seeds. After the second step, pruning reduces the number of candidates to the most promising $w$ beams, decreasing the complexity to $O(k\times n\times w)$.
To further reduce complexity, a heuristic based on text relevance can be used to predict the top $T$ most relevant past images. By restricting the search space to the latents of these top $T$ candidates, the model improves efficiency by prioritizing contextually relevant images, particularly in the early steps. While the initial three steps remain computationally demanding due to the expansive search space, pruning combined with this heuristic can significantly lower the computational cost by refining the candidate set.
Appendix F Application Examples
Images and other visual aids are essential tools that significantly improve comprehension in various fields. Whether while breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that makes the material easier to understand [26]. From technical procedures and instructional design to storytelling and educational content, the use of multiple-scene images allows the reader to grasp information more deeply and intuitively, transforming even the most complex or dense topics into something more accessible [39, 11]. However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence is a major challenge  [25].
Figure 15: Visualization of cross-attention in latent diffusion across different sets of latents. The heat maps reveal how attention patterns evolve at various stages, highlighting the alignment between text prompts and distinct latent representations.
F.1 Beam Search Tree
(a) The beam denoising search tree.
(b) Example of one complete decoded denoising process using contextual captions.
Figure 16: The beam diffusion model works in the denoising latent space by denoising the initial seed to latents and them decoding different beam paths. This shows how our method evolves in the latent space and how it can explore the latent space.
In Figure 16(a), we provide another example of how the latent search occurs within the beam denoising process. This illustrates the BeamDiffusion model at work, showing the denoising of the initial seed to latents and how different beam paths are decoded. The figure highlights how our method explores the latent space, iterating through various latents based on the beam search strategy.
In Figure 16(b), we observe that the generation of the selected image at step 4 behaves similarly to an isolated diffusion process. However, it is important to note that this process is influenced by multiple prompts. Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4. This demonstrates how the latent search process incorporates context from various stages, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
F.2 Sequence Examples
To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 20-28. These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 20, 20, 24, and 24. They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps. This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt. In contrast, GILL exhibits a different behavior. While it maintains high visual consistency, with all images in the sequence being visually similar, it fails to accurately follow the step-by-step textual instructions. GILL tends to generate images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete.
Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency. As shown in Figures 20, 20, 24, and 24, it generally produces the most coherent sequences compared to other methods. In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps. This combination of textual alignment and visual stability leads to more coherent and contextually accurate image sequences. While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches. However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 28, 28, and 28. The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence. While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
Figure 17: Example of the Recipes domain task ”Vegan Basil Pesto and Sun-Dried Tomato Pinwheels”, with a sequence of 4 steps.
Figure 18: Example of the Recipes domain task ”Tangy lemon sheet cake”, with a sequence of 4 steps.
Figure 19: Example of the Recipes domain task ”Zucchini, Pecorino & Mint Salad”, with a sequence of 4 steps.
Figure 20: Example of the Recipes domain task ”Bone Broth”, with a sequence of 3 steps.
Figure 21: Example of the out-of-domain VIST story ”Klamath Project Construction”, with a sequence of 5 steps.
Figure 22: Example of the out-of-domain VIST story ”Yosemite Weekend”, with a sequence of 5 steps.
Figure 23: Example of the out-of-domain VIST story ”Fireworks - 4th July in Washington, DC”, with a sequence of 5 steps.
Figure 24: Example of the out-of-domain VIST story ”Independence Day Parade”, with a sequence of 5 steps.
Figure 25: Example of the out-of-domain DIY task ”How to Clean Silk Rugs”, with a sequence of 2 steps.
Figure 26: Example of the out-of-domain DIY task ”How to Polish Wood”, with a sequence of 3 steps.
Figure 27: Example of the out-of-domain DIY task ”How to Preserve Basil”, with a sequence of 6 steps.
Figure 28: Example of the out-of-domain DIY task ”How to Clean Natural Stone”, with a sequence of 5 steps.

### v2

@@
 Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions.
-To address beam search’s quadratic complexity, we integrate a cross-attention mechanism that efficiently scores search paths and enables pruning, prioritizing alignment with both textual prompts and visual context.
-Human evaluations confirm that our approach outperforms baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment.
-By bridging advances in search optimization and latent space refinement, this work sets a new standard for structured image sequence generation.
+As the latent denoising space is explored, the beam search graph is pruned with a cross-attention mechanism that efficiently scores search paths, prioritizing alignment with both textual prompts and visual context.
+Human and automatic evaluations confirm that BeamDiffusion outperforms other baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment.
+All data and code implementations will be released after publication.
 1 Introduction
-Figure 1: BeamDiffusion models leverage search in the latent denoising space for the best sequence of denoised images, where $I_{n}$ represents the image at step $n$ and $S_{n}$ represents the text associated with step $n$.
 Image diffusion models [43, 9, 44] have made significant advancements in generating high-quality images.
 These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [35, 29, 38].
-While they excel at generating individual images based on specific text prompts, they face challenges when it comes to creating coherent sequences of images.
-In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across multiple steps.
+Despite the impressive results at generating individual images based on specific text prompts, diffusion models face challenges when it comes to creating coherent sequences of images.
+In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across a sequence of correlated prompts.
 This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence.
@@
 Beam search is a well-established technique in text generation tasks like machine translation [14, 51, 17, 37], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions.
-However, despite its success in other domains, it has been barely explored in our context, leaving its potential untapped.
-In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting text prompts or adding semantic cues.
+However, despite its success in other domains, it has been barely explored in this context, leaving its potential untapped.
+In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues.
 In this work, we explore beam search for latent prior selection in image sequence generation.
@@
 We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences.
-As illustrated in Figure 1, our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
-Due to the quadratic complexity of beam search algorithm, one needs to prune the search paths to keep a fixed number of candidate paths.
+As illustrated in Figure 1(a), our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
+As the beam diffusion algorithm traverses the latent space, the number of hypothesis grow and the least probably search paths are pruned to keep a fixed number of candidate paths.
 To this end, we incorporate a cross-attention mechanism [19] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps.
 This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
-We evaluated the effectiveness of our approach on three datasets: in the Recipes domain, and in the out-of-domain DIY Tasks and Visual Storytelling.
-Human evaluations show that our model outperformed baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
-2 Related Work
-Decoding methods are essential for enhancing the quality of large language models (LLMs).
-Techniques such as beam search, top-k and nucleus sampling are widely studied for their effectiveness in balancing computational cost with generating coherent, contextually appropriate outputs [40, 13, 9].
-Beam search has also proven effective in tasks like machine translation [27], image captioning [45, 19, 1], and visual storytelling [15].
-Building on these ideas, our work innovates by applying beam search to explore the latent space in diffusion models for image sequence generation.
-Ensuring coherence remains a significant challenge in multimodal sequence generation [10, 24, 6, 20].
-Some approaches, like AR-LDM [28], encode the context of caption-image pairs into multimodal representations to maintain consistency across outputs.
-Similarly, Rahman et al.
-[31] incorporates the history of U-net [35, 38] latent vectors to preserve narrative coherence.
-However, these methods often face challenges such as computational inefficiency and the risk of noise propagation over time.
-Other approaches, such as Gill [22], combine LLMs with image encoder-decoder systems.
-While effective for certain tasks, these models suffer from alignment issues, with retrieved or generated images failing to match the narrative context, and they too often require complex training.
-Our method avoids these pitfalls by direct conditional generation against the guiding text while explicitly maintaining coherence via the beam search.
-Interpreting diffusion models has become a focal point in recent research.
-Pont-Tuset et al.
-[30] anchor each word in textual descriptions to specific regions of images.
-Tang et al.
-[44] analyze text-image interactions through cross-attention maps in the denoising module, offering insights into how textual prompts influence the generated images.
-Dewan et al.
-[7] enhance this interpretability by leveraging partial information decomposition to analyze how various components of a text prompt impact the generated image.
-In contrast, our beam-search approach focuses on generating images directly from textual input.
-Rather than mapping text to pre-existing images, it generates image sequences that are visually coherent and are consistently aligned with the evolving sequence of textual narrative.
-Bordalo et al.
-[4] propose to use latent information from previous steps to guide coherent generation of subsequent images.
-However, their method relied on a greedy heuristic-based selection.
-CoSeD [32] improved on this idea and introduced a contrastive learning approach [5] to rank images generated from different latent seeds.
-Yet, both approaches focused on local refinements rather than optimizing across the entire sequence.
-Our beam search approach lifts these local selection criteria to perform global optimization.
-3 Synthesis of Non-Linear Visual Sequences
-In the following section, we define the challenges in image sequence generation, recall the core principles of LDM, Sec.
-3.1, and then we detail BeamDiffusion key aspects in terms of diffusion beams (Sec.
-3.2.1) and pruning (Sec.
-3.2.2) with a contrastive approach (Sec.
-3.2.3).
-3.1 Problem Formulation
-Generating a full sequence of visually coherent images from textual scene descriptions presents a challenge.
-Given a sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each scene $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps.
-Latent Diffusion Models (LDMs) [34] generate high-quality images in a compressed latent space.
-Instead of directly operating in pixel space, LDMs encode an input image $I$ into a latent representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [2].
-The U-Net backbone [35], denoted as $\epsilon_{\theta}$, incorporates a cross-attention mechanism [17] to condition the diffusion process on an external input $\tau_{\theta}(s)$, where $s$ represents a textual prompt.
-LDMs are trained by minimizing:
-$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right].$ (1)
-While effective, LDMs generate images independently, starting from a randomly initialized latent $z_{T}$.
-This lack of temporal coherence makes it challenging to ensure consistency across generated sequences, especially in non-linear narratives [41], where maintaining character identity, lighting, and environmental details are crucial.
+We evaluated the effectiveness of BeamDiffusion in three domains: the Recipes domain and in the out-of-domain DIY and Visual Storytelling.
+Results obtained with human annotations, automatic evaluations and LLM-as-a-judge, all showed that the proposed approach outperforms baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
 (a) The beam denoising search tree.
@@
 This shows how our method evolves in the latent space and how it can explore the latent space.
+2 Related Work
+Decoding methods are crucial for enhancing the quality of large language models (LLMs).
+Techniques such as beam search, top-k and nucleus sampling are widely studied for their effectiveness in balancing computational cost with generating coherent, contextually appropriate outputs [40, 10, 7].
+Beam search has also proven effective in tasks like machine translation [25], image captioning [48, 15], and visual storytelling [12].
+Ensuring coherence remains a major challenge in multimodal sequence generation [8, 22, 4, 16].
+Some approaches, like AR-LDM [28], encode the context of caption-image pairs into multimodal representations to maintain consistency across outputs.
+Likewise, Rahman et al.
+[32] incorporates the history of U-net [36, 39] latent vectors to preserve narrative coherence.
+Other approaches, such as Gill [18], combine LLMs with image encoder-decoder systems.
+StoryDiffusion [52], changes the U-Net’s cross-attention mechanism enabling training-free consistency by sampling tokens from previous images to maintain coherence in newly generated frames.
+StoryGen [21] denoises the previous image and the resulting features serve as conditioning for the subsequent generation of images.
+GenHowTo [45] uses a control net during inference to project the previous image directly into the latent space to guide the diffusion process for the next image step.
+StackedDiffusion [26] processes stacked image sequences as a single latent representation allowing the attention mechanism to consider all steps together, enhancing coherence across steps.
+In [2], authors proposed to use latent information from previous steps to guide coherent generation of subsequent images relying on a greedy heuristic-based selection.
+CoSeD [33] improved on this idea and introduced a contrastive learning approach [3] to rank images generated from different latent seeds.
+However, both approaches focused on local refinements rather than optimizing throughout the sequence.
+Our beam search approach extends these local selection criteria to perform global optimization.
+While effective for certain tasks, these models suffer from alignment issues, with retrieved or generated images failing to match the narrative context, and they too often require complex training.
+Our method addresses these previous constraints by generating images directly from the guiding text, maintaining consistency via a beam search for optimal seeds, similar to methods such as [41] that view image generation as a search process.
+3 Generating Non-Linear Visual Sequences
+Generating a full sequence of visually coherent images from textual scene descriptions is a challenging research problem.
+Given a sequence of steps $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each sequence step $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps of the sequence.
+While Latent Diffusion Models (LDMs) [35] can generate high-quality images in a compressed latent space, this latent space is designed to generate images independently, starting from a randomly initialized latent $z_{T}$.
+LDMs encode an input image $I$ into thelatent space representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [35].
+The reverse diffusion process is modeled by a U-Net backbone [36], $\epsilon_{\theta}$, which incorporates a cross-attention mechanism [47] to condition the denoising iterations on an external input, e.g.
+the output embeddings of a text encoder $\tau_{\theta}(s)$.
+LDMs are trained to iteratively estimate a denoised variant of $z$, with the loss
+$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
+This loss evidences the lack of awareness of other denoising processes, making it challenging to ensure consistency acrosscorrelated prompts or previous denoising processes.
+This is particularly challenging in non-linear narratives [42], where maintaining character identity, lighting, and environmental details are crucial.
+To address these challenges, prior work [32, 21, 45, 2, 33] reuse latent denoising iterations from earlier steps in the sequence, paving the way to a key question: how to determine the optimal set of denoising latents that maximizes coherence across different generations?
+The answer lies in a combinatorial search space, that BeamDiffusion can traverse with a search strategy to determine an optimal solution, as we will discuss next.
 3.1 Beam Diffusion Model
@@
 The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space.
-BeamDiffusion incorporates latent representations $\mathcal{Z}_{j}$ from previously generated images.
+BeamDiffusion incorporates latent representations from previously generated images.
 These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence.
-This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the new scene description.
-In the following sections, we will detail the proposed method.
+This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the next scene description.
+In the remainder of this article, the term step will always refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
 3.1.1 Latent Space Exploration with Diffusion Beams
-To maintain both diversity and coherence in the generated image sequence, we propose a diffusion-based exploration of the latent space across multiple steps.
-At each generation step $j$, the set $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots,I_{j}^{(B)}\}$ consists of all possible candidate images, with $I_{j}^{(b)}$ representing the image generated at step $j$ for beam $b$.
+To generate coherent and diverse image sequences, we explore the latent space of diffusion models using a beam search-inspired strategy.
+For each sequence step $j$ we generate a set of candidate images $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots\}$, where each $I_{j}^{(k)}$ is conditioned on a different beam path.
+Algorithm 1 BeamDiffusion Algorithm
+1:  Input parameters: $\{s_{1},\ldots,s_{N}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
+2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$triangleright Beam initialization with $r$ different images.
+3:  for $j=2$ to $L$ do
+4:     for all beam $b\in\hat{B}_{j-1}$ do
+5:        $\mathcal{Z}_{j}\leftarrow$$\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$triangleright Read denoising latents from previous steps.
+6:        for all latent $z\in\mathcal{Z}_{j}$ do
+7:           $I_{j}\leftarrow\text{SD}(s_{j},z)$triangleright Generate candidate a image.
+8:           $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$triangleright Extend beam with candidate image.
+9:        end for
+10:     end for
+11:     $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$triangleright Score the new image of the beam.
+12:     $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$triangleright Prune beams and keep top $w$ beams.
+13:  end for
+14:  return  best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
 Sampling the Denoising Latent Space.
-In the first step of the sequence, $s_{1}$, the model generates multiple candidate images using different random seeds.
-This approach allows for a broader exploration of the latent space and reduces Bayes risk [3], as described in Appendix B.
-Hence, in this initial step, $s_{1}$, we obtain the first exploration samples from the denoising latent space, denoted as $\mathcal{Z}_{1}=\{z_{1}^{(1)},\ldots,z_{1}^{(K)}\}$.
-Building on prior work [4, 32], incorporating latent denoising from earlier steps ($s_{<j}$) can enhance the denoising process.
-We define $\mathcal{Z}_{j}$ to be the accumulation of latent representations from all preceding images up to step $j$:
-$\mathcal{Z}_{j}=\bigcup_{i\in\{j-n,\ldots,j-1\}}\{z_{i}^{(1)},\dots,z_{i}^{(K)}\},$ (2)
-where each latent representation $z_{i}^{(k)}$ corresponds to the $k$-th iteration of the diffusion process for image $I_{i}$.
-By utilizing latents from $n$ previous stages (steps back), our approach ensures contextual consistency across generations.
-Considering one particular search beam $b$ in the latent space, to generate a new candidate image $I_{j}^{(b)}$, we condition the diffusion process on a latent representation $z$ selected from the accumulated latents in $\mathcal{Z}_{j}$, as represented in Figure 2(a).
-This conditional dependency is captured by the following probability expression:
-$\small p\left(I_{j}^{(b)}\mid\theta,z\in\mathcal{Z}_{j},s_{j}\right),$ (3)
-where $\theta$ are the model parameters.
-This formulation ensures that all denoising processes are correlated, as each new image $I_{j}^{(b)}$ is generated based on latents $z$ drawn from $\mathcal{Z}_{j}$.
-In this way, each iteration in the denoising process acts as a seed for the next, establishing a dependency between successive denoising processes (Figure 2(b)).
+In our approach, the denoising process for each sequence step is conditioned by representations from prior steps, helping the model maintain coherence across the sequence.
+In the first step of the sequence, $s_{1}$, we generate multiple candidate images by running the diffusion process with $r$ random noise seeds (Alg.
+1, line 2).
+This strategy enables broader exploration of the latent space and reduces Bayes risk [1], as detailed in Appendix D.
+For every generation process corresponding to a sequence step $s_{i}$, we store the first $N$ latent space vectors $z_{i}=\{z_{i,t}\}_{t\in\subseteq\{1,\ldots,N\}}$ produced during the denoising iterations.
+On each step prompt $s_{j>1}$ we need to generate diverse samples that (a) explore different denoising trajectories in the latent space, and (b) are strongly correlated with previous samples.
+Leveraging the stored latent space vectors, we sample the latent subspace of $n$ previous denoising processes.
+Hence, for steps $j>1$, to enable contextual consistency, we first accumulate representations from the previous $m$ steps:
+$\mathcal{Z}_{j}=\bigcup_{i\in\{j-m,\ldots,j-1\}}\{z_{i}\}.$ (2)
+To generate new candidate images $\mathcal{I}_{j}=\{I_{j}^{(1)},\ldots,I_{j}^{(k)},\ldots,I_{j}^{(N\cdot m)}\}$ at step $j$, we condition the diffusion model on the latent trajectories in $\mathcal{Z}_{j}$.
+Specifically, for each latent trajectory $z\in\mathcal{Z}_{j}$, the model performs a new diffusion process to produce a corresponding image candidate $I_{j}^{(k)}$, Alg.
+1, line 5 and 7.
+This step establishes a dependency between generations, where latents from earlier steps guide the synthesis of visually coherent transitions over time, as illustrated in Figure 1.
 Diffusion Beams.
-Following the idea of beam search, which maintains a set of the most promising candidates at each step and extends them sequentially [46], we construct diffusion beams by iteratively expanding candidate images.
-Formally, for each step $j$, we generate the set of all candidate images $\mathcal{I}_{j}$ to compute the set of candidate diffusion beams $\hat{B}_{j}$.
-At the initial step $j=1$, there are no previous beams to combine, so the initial set of beams can be represented as,
-$\small\hat{B}_{1}=\bigcup_{k}\{I_{1}^{(k)}\mid I_{1}^{(k)}\in\mathcal{I}_{1}\},$ (4)
-where $\mathcal{I}_{1}$ is the set of initial candidate images generated using different random seeds indexed by $k$.
-For subsequent steps $j>1$, each beam $\hat{B}_{j-1}^{(b)}$ from the previous step is extended with new candidate images $I_{j}^{(b\cdot k)}$ from the set $\mathcal{I}_{j}$, and this extension is represented by the operator $\oplus$.
-The new set of beams is now represented as,
-$\small\hat{B}_{j}=\bigcup_{(b,k)}\left\{\hat{B}_{j-1}^{(b)}\oplus I_{j}^{(b\cdot k)}\mid I_{j}^{(b\cdot k)}\in\mathcal{I}_{j}\right\}$ (5)
-where $b$ is the index of a beam from the set $\hat{B}_{j-1}$, and $k$ is the $k$-th latent variable from $\mathcal{Z}_{j}$.
-For each combination of a previous beam $\hat{B}_{j-1}^{(b)}$ and latent $\mathcal{Z}_{j}^{(k)}$, the model considers a new candidate image $I_{j}^{(b\cdot k)}$ from the set of candidate images $\mathcal{I}_{j}$.
-The union over $(b,k)$ combines all possible sequences of previous beams with the new candidate images, expanding the search space for the next step of generation.
+Inspired by beam search [50], which maintains and extends the most promising candidates at each sequence step, we introduce diffusion beams to structure sequential image generation.
+At each step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step $j-1$ and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$.
+The resulting beams are defined as
+$\hat{B}_{j}=\bigcup_{(b,k)}\left\{\hat{B}_{j-1}^{(b)}\oplus I_{j}^{(b\cdot k)}\mid I_{j}^{(b\cdot k)}\in\mathcal{I}_{j}\right\},$ (3)
+where $\oplus$ denotes the concatenation of a prior beam with a new candidate image, and $(b\cdot k)$ indexes the pairing of beam $b$ and candidate image $I_{j}^{(k)}$.
+This process builds a tree of candidate sequences, Alg.
+1, line 8, enabling the model to explore diverse yet coherent visual narratives over time.
 3.1.2 Pruning Diffusion Beam Paths
-Tracking diffusion beams helps maintain context and coherence across the growing image sequence by exploring multiple potential continuations.
-However, as the sequence progresses, focusing on the most promising paths within the beam budget (the number of beams that can be maintained simultaneously) becomes increasingly important to ensure the coherence of the generated images and optimize computational resources.
-To achieve this, our model prunes the diffusion beam paths, narrowing down the search space for the best candidates while promoting diversity and consistency in the maintained beams.
-To prune candidate beams, the model scores each candidate $\hat{B}_{j}$ by the likelihood of image $I_{j}$ in the beam given the previous images $I_{<j}$ and the step text.
-This is represented as $p(I_{j}|s_{j},B_{j-1}^{(b)})$, where $B_{j-1}$ represents the set of the most promising candidate beams from the previous step.
-Formally, the scoring is defined as:
-$\small score(\hat{B}_{j}^{(b)})=\sum_{i=1}^{j}\log p(I_{i}|I_{1},\dots,I_{i-1},s_{1},\ldots,s_{i}),$ (6)
-where we sum the individual log-probabilities of each image in the sequence, reflecting the overall coherence and alignment with the context set by the previous images and prompts.
-After scoring all candidate images in $\mathcal{I}_{j}$, the search space is refined by performing a beam search operation.
-This operation selects the most promising $w$ diffusion beams, effectively limiting the beam width to $w$.
-The refinement process is expressed as,
-$\small B_{j}=\mathop{\mathrm{argmax}}_{B_{j}\subseteq\hat{B}_{j},|B_{j}|=w}\left(\sum_{i=1}^{w}score(\hat{B}_{j}^{(i)})\right)$ (7)
-where, $B_{j}$ represents the most promising $w$ candidate diffusion beams at the step $j$, ensuring that the updated beam evolves to explore the most promising options, guided by the previous context and the newly generated candidates, while maintaining a controlled beam width.
-To maintain a broad search space in the early stages, pruning is delayed until after the second step.
-During this initial phase, all candidates are retained for exploration.
-After the second step, pruning is applied, limiting the candidates to the $w$ most promising.
-This approach balances exploration and computational efficiency, focusing the model’s efforts on the most promising candidates in the later stages of generation.
-Figure 3: Visualization of cross-attention in latent diffusion, showing how BeamDiffusion focuses on different regions of the image based on the prompt at each stage.
-The heatmap highlights the areas with the strongest attention.
-3.2.3 Contrastive Score
-While the basic pruning mechanism outlined above provides a foundation for selecting candidate images based on raw probabilities, the scoring function is now implemented as contrastive classifier.
-This classifier improves the image selection by learning from past sequences and aligning the candidates with the evolving context [32].
-As illustrated in Figure 3, the heat maps show how BeamDiffusion attends to different regions of the input text at various stages, highlighting the evolving alignment between textual prompts and image features.
-To compute these probabilities, we refine the approach using a softmax-based method to determine the likelihood of each candidate image $I_{j}$ relative to all other possible candidates in the sequence.
-In this refined process, the probability $p(I_{j}|s_{j},B_{j-1}^{(b)})$ is replaced by the output of the classifier $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$, which approximates the likelihood of each candidate fitting the sequence.
-The more detailed version of the scoring mechanism can be expressed as,
-$\scriptsize\begin{split}score(\hat{B}_{j}^{(b)})&=\sum_{i=1}^{j}\log\left(\frac{\exp\left(\varphi(I_{i},s_{i},B_{i-1}^{(b)})\right)}{\sum_{I_{i}^{\prime}\in\mathcal{I}_{i}}\exp\left(\varphi(I_{i}^{\prime},s_{i},B_{i-1}^{(b)})\right)}\right),\end{split}$ (8)
-where the denominator in the softmax function normalizes the scores by summing over the candidate images $I_{i}^{\prime}$ for beam $b$ within the set $\mathcal{I}_{i}$.
-This ensures that the score of selecting a particular candidate image $I_{j}$ is computed relative to all other candidate images in the sequence, capturing how well it fits the overall sequence in terms of visual content and textual alignment.
+Tracking multiple diffusion beams enables BeamDiffusion to maintain visual coherence while exploring diverse continuations across a sequence.
+However, as the sequence length increases, the number of candidate beams grows rapidly, making it computationally infeasible to retain them all.
+To address this, we apply a pruning strategy based on a learned contrastive scoring function, which allows us to retain only the most promising beams while preserving diversity.
+At each sequence step $j$, we construct a set of candidate beams $\hat{B}_{j}$.
+Each candidate beam $\hat{B}_{j}^{(b)}$ is formed by extending a beam from the previous step $\hat{B}_{j-1}^{(b)}$ with a newly generated image $I_{j}^{(k)}$.
+A complete candidate beam at step $j$ is a sequence of images $\hat{B}_{j}^{(b)}=[I_{1}^{(b)},I_{2}^{(b)},\ldots,I_{j}^{(b)}]$, one per step.
+To evaluate the overall quality of a beam, we use a contrastive classifier [33], denoted by $\varphi$.
+Specifically, for each image $I_{i}^{(b)}$ in the beam, the classifier $\varphi$ computes a compatibility score based on the image’s prompt $s_{i}$ and the previous images in the beam $B_{i-1}^{(b)}$, which includes earlier images and prompts in the same beam.
+The total beam score is then the sum of these per-image scores over all sequence steps up to $j$:
+$score(\hat{B}_{j}^{(b)})=\sum_{i=1}^{j}\log\left(\frac{\exp\left(\varphi(I_{i}^{(b)},s_{i},B_{i-1}^{(b)})\right)}{\sum_{I_{i}^{\prime}\in\mathcal{I}_{i}}\exp\left(\varphi(I_{i}^{\prime},s_{i},B_{i-1}^{(b)})\right)}\right),$ (4)
+where $\mathcal{I}_{i}$ is the set of all candidate images at sequence step $i$.
+The contrastive formulation ensures that each image is evaluated not in isolation, but relative to other candidates, capturing how well it fits the evolving sequence context.
+Once scored, we prune the beam candidates by selecting the top $w$ scoring beams:
+$B_{j}=\mathop{\mathrm{argmax}}_{B_{j}\subseteq\hat{B}_{j},|B_{j}|=w}\left(\sum_{i=1}^{w}score(\hat{B}_{j}^{(i)})\right),$ (5)
+where $B_{j}$ denotes the final set of retained beams for sequence step $j$.
+To balance exploration and efficiency, pruning is disabled for the initial steps, allowing all candidates to be retained initially.
+After a few step, we apply pruning to keep only the top $w$ beams.
+This strategy ensures wide exploration early on and focused, high-quality continuation in later sequence steps.
 Training.
-To train the classifier effectively, we employ a contextual sequence training approach, which enables the model to capture sequential relationships across different tasks.
-During training, the true label $l_{t,j}$ indicates whether the candidate image $I_{j}$ fits correctly into the sequence, given the previous steps $B_{j-1}^{(b)}$ of the $b$-th beam.
-Here, $t$ refers to the task index and $j$ represents the step index within the sequence.
-The model’s prediction $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$ is based on the candidate image $I_{j}$, its textual description $s_{j}$, and the context provided by the prior steps.
-The model aims to minimize cross-entropy loss by aligning the candidate image with its context, ensuring it fits well within the sequence according to the true labels $l_{t,j}$, defined as:
-$\small\operatorname*{argmin}\sum_{t=1}^{M}\sum_{j=1}^{N}l_{t,j}\log(\varphi(I_{j},s_{j},B_{j-1}^{(b)}))$ (9)
-Through contrastive training, the model learns to align representations of similar steps while enhancing the distinction between steps that differ semantically or visually.
-3.2.4 Contextual Scene Descriptions
-In many cases, step descriptions alone may lack context as precise visual prompts.
-These step descriptions are often interdependent, relying on prior steps for contextual understanding, or may miss visual details, e.g.
-”Put the ice cream in freezer for 1 hour”, or actions that involve multiple tasks, such as ”Slice the tomatoes, grate the cheese, and stir the sauce”.
-To address this challenge, we leverage Gemini [33] to refine step descriptions into visually detailed prompts while considering sequence context.
-Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
-$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (10)
-Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps. For further information, see Appendix E.
-Recipes DIY VIST
-Method Human (%) Gemini (%) Human (%) Gemini (%) Human (%) Gemini (%)
-Greedy/CoSeDlen=1 [32] 13.3 16.7 20.0 10.0 27.2 16.7
-Nucleus Sampling [14] 33.3 16.7 30.0 30.0 18.2 26.6
-GILL [22] 0.0 3.3 10.0 20.0 0.00 0.0
-BeamDiffusion 53.3 63.3 40.0 40.0 54.5 56.7
-Table 1: Frequency of method selection by human annotators and Gemini.
+The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well align an image is with the previously sequence of steps and images.
+For a given sequence $t$ and sequence step $j$, the ground-truth label $l_{t,j}$ indicates whether image $I_{j}$ belongs to the sequence given its prompt $s_{j}$ and context $B_{j-1}^{(b)}$.
+Negative examples are hallucinated from other sequences.
+The model minimizes the cross-entropy loss encouraging the model to promote candidates that best match the sequence’s semantic and visual dynamics.
+See appendix A for implementation details.
 4 Experimental Setup
 In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences.
-We outline the datasets used, identify other baseline methods, and explain the human and Gemini evaluations focused on semantic and visual consistency across generated sequences.
-All data and implementations will be released after publication.
+We outline the datasets used, identify other baseline methods, and explain the human, automatic and LLM-as-a-judge (Gemini [34]) evaluations focused on semantic and visual consistency across generated sequences.
 Datasets.
@@
 These out-of-domain datasets allow us to assess the generalizability of our model across different task domains.
-More details can be found in Appendix D.
+More details can be found in Appendix B.2.
 Baselines.
-To assess the performance of our image sequence generation method for multistep tasks, we compared BeamDiffusion to two other decoding approaches, as well as GILL [22], focusing on different strategies for generating consistent image sequences:
-  • Greedy/CoSeDlen=1 [32]: In this method, a new image is generated at each step using the first five latents from the previous image, selecting the best overall image.
-Although similar to CoSeDlen=1, it relies solely on CLIP without alterations, whereas CoSeD uses a trained module for selection.
-  • Nucleus Sampling [14]: Rather than choosing the single most likely image, nucleus sampling randomly selects from the set of images whose cumulative probability exceeds a specified threshold $p$, according to CLIP.
-This allows for more diversity in the generated sequence while still emphasizing high-probability outcomes.
-  • GILL [22]: Unlike the other methods, GILL does not explicitly search the latent space.
-Instead, it ensures coherent image sequences by iteratively generating images, while maintaining consistency across the sequence.
-GILL achieves this by deciding when to retrieve an image from a candidate set or when to generate a new image, based on the context of the previous images and text inputs.
+To evaluate the performance of our image sequence generation method, we compared BeamDiffusion with SD2.1 [35] against two other competing approaches to generate image sequences: GILL [18] and StackedDiffusion [26].
+Additionally, we considered two decoding strategies: Greedy and Nucleus Sampling [11].
+As these decoding strategies are novel in the context of image generation, we provide a brief explanation of each.
+All methods used 50 denoising iterations.
+In the Greedy approach, a new image is generated at each step using the first five latents from the previous image, with the image that scores highest overall being selected.
+Although this is similar to CoSeDlen=1 [33], it differs in that it relies exclusively on CLIP for selection without any learned module, whereas CoSeD incorporates a trained selection mechanism.
+In contrast, Nucleus Sampling introduces diversity by sampling from the subset of images whose cumulative CLIP-assigned probability exceeds a predefined threshold $p$.
+Rather than always selecting the most likely image, this method favors high-probability outcomes while allowing for variation in the generated sequence.
 Human Evaluation.
-In the first task, the annotators compared pairs of beam search configurations in a ”battle-style” format to select which beam configuration was the best regarding beam width ($w$ in 7) and steps back ($n$ in 7).
+To find the optimal hyperparameters, the annotators compared pairs of beam search configurations to select the optimal beam width ($w$ in Eq.
+5) and steps back ($m$ in Eq.
+2).
 The winning configuration is compared against other configurations until a new configuration is selected as better.
@@
 Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
-Having chosen the optimal configuration, we need to assess how it compares to other methods.
-In the second annotation task, BeamDiffusion was compared with our baseline models.
-The annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency, with the possibility of choosing multiple options or none.
-For further information, see Appendix C.
+Having chosen the optimal configuration, we compared BeamDiffusion to other methods.
+In the second annotation experiment, BeamDiffusion was compared with other baseline models.
+Annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency with the possibility of choosing multiple options or none.
+For further information, see Appendix B.1.
 Gemini Evaluation.
 To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines.
-Its selection process closely mirrored that of human annotators, with one key difference: Gemini tends to default to ”both good” when given the option.
-To avoid this, we removed the ”both good” and ”both bad” options from its evaluation to ensure more accurate results.
-We also refined our method for determining the best generation technique.
-Initially, presenting all four sequences at once caused the model to focus primarily on the last two, leading to biased results.
-To address this, we switched to pairwise evaluations, comparing sequences two at a time, which provided a more balanced and reliable assessment.
-For details, please refer to Appendix C.
+Its selection process closely mirrored that of human annotators.
+For details, please refer to Appendix B.1.
 Human-Gemini Agreement.
-We conducted an alignment assessment between human annotators and Gemini.
-We asked 10 annotators to rate 16 Recipes, 10 DIYs, and 10 VIST tasks.
-After this assessment, we computed Fleiss’ kappa [8], obtaining a value of 70%, indicating a substantial agreement.
-Given this substantial alignment, we extended the evaluation to the remaining sequences with Gemini.
+We conducted an alignment assessment to measure the agreement between human annotators and Gemini.
+For this, human annotators were asked to rate 16 Recipes, 10 DIYs, and 10 VIST tasks.
+These ratings were then compared to the evaluations made by Gemini.
+To quantify the agreement between the two, we computed Fleiss’ kappa [6], obtaining a value of 70%, which indicates a substantial level of agreement.
+Based on this strong alignment, we extended the evaluation to the remaining sequences using Gemini.
+Automatic Metrics.
+Following prior work [2, 33, 21], we evaluate each method using: (1) Visual Consistency, based on average image similarity via DINO [27] (DINO-I) and CLIP [31] (CLIP-I); and (2) Semantic Alignment, using CLIP (CLIP-T) to measure similarity between the current image and the average of all step texts up to and including the current one.
+To place the scores on a comparable scale, we normalize CLIP-I, CLIP-T and DINO by their maximum.
+Finally, we compute CLIP* as the product between the CLIP-I and CLIP-T, and DINO* as the product between the DINO-I and CLIP-T, this formulations balances visual and semantic quality, while emphasizing semantic alignment, which is critical for accurately reflecting the stepwise input text.
+See Appendix B.1 for details.
+5 Results and Discussion
+In this section, we describe the performance of BeamDiffusion based on both human and automatic evaluations.
+We begin by analyzing the human evaluation, proceed to discussing the automatic evaluations, then focus on specifics of our approach such as hyperparameters and non-linear decoding characteristics, and finally provide a qualitative analysis.
+Additional statistical information about the results can be found in Appendix B.1.
+5.1 General Results
+Table 1 compares all methods across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain).
+BeamDiffusion consistently outperforms other methods.
+In the context of Recipes, BeamDiffusion is selected 50.0% of the time by humans and 37.6% by the LLM, with Nucleus and Greedy following.
+In out-of-domain areas, it is chosen once more 50.0% of the time by humans and 30.0% by the LLM-judge for the DIY domain, and 58.3% for humans and 34.9% by the LLM for the VIST dataset.
+Among sequence generation methods, GILL consistently underperforms across all domains, receiving at most 3.3% of human votes and 16.1% from the LLM in VIST.
+StackedDiffusion performs better than both Greedy and GILL, but still lags behind BeamDiffusion, with 20.0% of human selections and 9.8% of LLM-judge selections in the Recipes domain.
+Overall, BeamDiffusion excels in Recipes and VIST, however performance in DIY remains challenging, with smaller gaps between methods due to the domain’s inherent complexity.
+Table 1: Frequency of method selection by human annotators(H) and Gemini(G).
+Recipes DIY VIST
+Method H (%) G (%) H (%) G (%) H (%) G (%)
+GILL [18] 0.0 2.0 0.0 20.0 3.3 16.1
+StackedDiffusion [26] 20.0 9.8 12.5 15.6 0.00 3.2
+Image Sequence Decoding
+Greedy/CoSeDlen=1 [33] 0 27.6 0 23.3 0 26.3
+Nucleus Sampling [11] 30.0 23.0 37.5 27.7 8.3 26.6
+BeamDiffusion 50.0 37.6 50.0 30.0 58.3 34.9
+5.2 Automatic Metrics
+Table 2 shows that BeamDiffusion achieves the highest CLIP* and DINO* scores (0.40 and and 0.30), demonstrating a strong balance between image consistency and semantic alignment.
+BeamDiffusion excels in maintaining a good compromise between text grounding and image similarity achieving the best result in image consistency (CLIP-I of 0.83 and DINO-I of 0.62).
+Similarly, for CLIP-T, BeamDiffusion achieves the highest score for image-text alignment (CLIP-T of 0.48).
+However, in this case, the highest result is shared, as Greedy and Nucleus obtain the same score.
+This outcome is expected because Greedy and Nucleus utilize CLIP-T to select their outputs by aligning the generated sequence with the text.
+In terms of sequence generation methods, StackedDiffusion performs better than GILL in all metrics but still lags behind BeamDiffusion and the other decoding methods.
+Table 2: Comparison of CLIP-I, DINO-I and CLIP-T across different methods
+Method CLIP-I DINO-I CLIP-T CLIP* DINO*
+GILL 0.55 0.49 0.43 0.23 0.21
+StackedDiffusion 0.76 0.46 0.47 0.36 0.22
+Image Sequence Decoding
+Greedy 0.81 0.54 0.48 0.39 0.26
+Nucleus 0.77 0.53 0.48 0.37 0.25
+BeamDiffusion 0.83 0.62 0.48 0.40 0.30
+5.3 Best Beam Search Configuration
+The hyperparameters beam width ($w$ in Eq.
+5) and steps back ($m$ in Eq.
+2), play a crucial role in the performance of BeamDiffusion.
+In this section, we report an exaustive evaluation on how different beam widths (2, 3, 4, and 6) and steps back (2, 3, and 4) improve the latent selection and the overall quality of the generation.
+A more detailed discussion on the impact of the latents selection is provided in Appendix C.
+We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix B.1 for details.
+To determine the best configuration, annotations were used to discard non-optimal configurations.
+The results in Table 4 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%.
+Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration.
+The configurations with beam width 6 performed the worst in all the settings tested.
+To assess the reliability of these findings, we used Gemini for the same evaluation on six tasks, achieving an 85% Fleiss’ Kappa agreement [6], indicating near-perfect alignment.
+The results from the extended Gemini evaluation (Table 4) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained the best, now achieving a 48.57% selection rate, a significant improvement over the 33.33% from the initial six-task evaluation.
+Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative.
+However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
+Table 3: Influence of different hyperparameter values according to human annotations.
 Steps Back Beam Width
@@
 4 16.67 8.33 8.33 -
-Table 2: Influence of different hyperparameter values according to human annotations.
+Table 4: Influence of different hyperparameter values according to Gemini annotations.
 Steps Back Beam Width
@@
 4 - - 2.86 -
-Table 3: Influence of different hyperparameter values according to Gemini annotations.
+5.4 Non-Linear Sequence Denoising
+To analyze how BeamDiffusion explores the latent space, we measured the number of times that a sequence step provides latents for the generation of subsequent steps of the sequence.
+In this experiment we consider sequences of four steps, (Figure 2).
+Appendix C.2 provides an analysis at the level of the denoising iterations.
+Latents from step 1 were chosen 69.9% of
 Figure 2: Non-linear steps selection, according to selection opportunities.
-5 Results and Discussion
-In the following sections, we demonstrate the performance of BeamDiffusion and examine different aspects: hyperparameters, non-linear decoding properties and a qualitative analysis.
-5.1 General Results
-The results in Table 1 compare the performance of four methods, Greedy/CoSeDlen=1, Nucleus Sampling, GILL, and BeamDiffusion, evaluated through human annotations and the Gemini model across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain).
-Across domains, BeamDiffusion consistently outperforms other methods.
-In the Recipes domain, it leads with 53.3% in human evaluation and 63.3% in Gemini.
-Nucleus follows with 33.3% (human) and 16.7% (Gemini), while greedy scores 13.3% (human) and 16.7% (Gemini).
-In DIY, BeamDiffusion achieves 40.0% in both evaluations, again outperforming greedy and nucleus.
-In VIST, BeamDiffusion leads with 54.5% (human) and 56.7% (Gemini), with greedy and nucleus scoring significantly lower.
-GILL consistently underperforms, receiving 0% selection in VIST across both evaluations.
-In Recipes, it scores just 3.3% in Gemini and remains the weakest performer across all methods.
-While it performs slightly better in DIY, it still lags behind other approaches.
-Although BeamDiffusion excels in Recipes and VIST, performance in DIY remains more challenging, with smaller performance gaps between methods.
-This reflects the inherent complexity of the DIY domain, where generating coherent outputs is particularly difficult.
-Overall, BeamDiffusion stands out as the most effective method across all domains, particularly in Recipes and VIST.
-However, its performance in DIY highlights the need for a more effective search in the latent space to ensure consistency and coherence across diverse tasks.
-Figure 5: Example of the Recipes domain task ”Jamaican Callaloo With Shrimp”, with a sequence of 4 steps.
-5.2 Best Beam Search Configuration
-The hyperparameters beam width ($w$ in Eq.
-7) and steps back ($n$ in Eq.
-2), play a crucial role in the performance of BeamDiffusion.
-In this section, we evaluated different configurations across four beam widths (2, 3, 4, and 6) and three steps back (2, 3, and 4).
-In this study, we specifically focus on latents indexes 0, 1, 2, and 3, denoted as $K$ in Eq.
-2, which yield the most promising results.
-A more detailed discussion on the impact of latents is provided in Appendix A.1.
-We first performed human annotations on six tasks to identify the optimal beam search configuration, see Appendix C for details.
-The results in Table 3 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%.
-Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed competitive performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration.
-The configurations with beam width 6 performed the worst in all the settings tested.
-To assess the reliability of these findings, we used Gemini for the same evaluation on six tasks, achieving an 85% Fleiss’ Kappa agreement [8], indicating near-perfect alignment.
-The results from the extended Gemini evaluation (Table 3) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained the best, now achieving a 48.57% selection rate, a significant improvement over the 33.33% from the initial six-task evaluation.
-Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative.
-However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
-5.3 Non-Linear Sequence Denoising
-To analyze how beam diffusion explore the latent space as formulated in Sec.
-3.2.1, we examined the step from which the latent representation was selected during the generation process.
-For an analysis of latent selection, see Appendix A.2.
-We focused the step selection process in four-step sequences (Figure 4).
-Step 1 was chosen 69.9% of the time, highlighting its key role in shaping the sequence.
+the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence.
 Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%.
-The sum of these percentages exceeds 100% because they represent the relative selection frequency of each step, considering the number of possible selection opportunities.
 Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate.
 Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total.
-This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate stages.
-By balancing local accuracy with global consistency, BeamDiffusion ensures high-quality image sequences that remain visually and contextually aligned.
-Figure 6: Example of the out-of-domain DIY task ”How to Make a Door”, with a sequence of 5 steps.
-Figure 7: Example of the out-of-domain VIST story ”July 4th Barbecue”, with a sequence of 5 steps.
+This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate sequence steps.
+(a) Recipes domain task.
+(b) DIY task.
+(c) VIST story.
+Figure 3: Example tasks from three different domains: Recipes ("Jamaican Callaloo With Shrimp"), DIY:("How to Make a Door"), and VIST ("July 4th Barbecue").
 5.5 Qualitative Analysis
-To illustrate how different decoding methods impact the semantic and visual consistency of generated output, we present several examples in Figures 57, and in Appendix F.2.
-In Figure 5, we see that BeamDiffusion consistently maintains the same pan across all steps, a quality not always achieved by greedy or nucleus methods.
-Meanwhile, GILL struggles to follow the text from step to step, leading to more inconsistencies in the outputs.
-Looking at the out-of-domain DIY and VIST datasets, in Figure 6, we observe a task with intricate details.
-Although other methods fail to consistently capture the color of the plywood, BeamDiffusion outperforms by maintaining this detail, demonstrating its strength in preserving fine-grained elements in complex tasks.
-In Figure 7, we again observe significant inconsistencies with the greedy and nucleus methods, which alternate between black-and-white and color images.
-In contrast, BeamDiffusion keeps the color consistent throughout the sequence.
-Additionally, background elements, such as the bushes, remain stable across the first three steps.
-Even the persons in first step closely resemble those in the third, further emphasizing the BeamDiffusion consistency.
-At the same time, GILL consistently struggles to follow the text, failing to capture the necessary details.
-This analysis highlights BeamDiffusion ability to maintain high levels of semantic and visual consistency, even in out-of-domain tasks, much like its performance on Recipes.
+To highlight how different decoding methods and models impact the semantic and visual consistency of generated outputs, we present several examples in Figures 3(a)3(c) and Appendix F.2.
+We compare three decoding methods (Greedy, Nucleus, and Beam) with two models (GILL and StackedDiffusion).
+In Figure 3(a), BeamDiffusion consistently maintains the same pan across all steps, a feature not always seen in Greedy or Nucleus methods.
+While GILL struggles to follow the step text and results in more inconsistencies, Beam performs better in preserving visual coherence compared to Greedy and Nucleus, though it still faces some challenges.
+For DIY tasks,as shown in Figure 3(b), BeamDiffusion can preserve basic elements such as the color of the plywood and a basic adherence to the text.
+However, overall, all methods face challenges in capturing precise details.
+In the VIST dataset, Figure 3(c), Greedy and Nucleus alternate between black-and-white and color images, while BeamDiffusion ensures color consistency throughout the sequence.
+Furthermore, background elements, such as bushes, are maintained in the first three steps, and the individuals in the first and third steps closely resemble each other.
+Overall, BeamDiffusion excels in maintaining both semantic and visual consistency, even in complex out-of-domain tasks, In contrast, methods such as StackedDiffusion and GILL encounter challenges in keeping a balance between text alignment and consistency across images.
+6 Limitations
+Exploring the latent denoising space with BeamDiffusion requires Diffusion Models that provide access to fine-grain latents across denoising iterations, where BeamDiffusion still has the opportunity to steer the reverse diffusion process.
+The theoretical complexity of BeamDiffusion is linear with the number of beams that we configure the model to explore.
+The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process.
+However, this provides a solid ground to explore exciting new ideas where stronger constraints can be enforced in the explored beams.
 7 Conclusions
@@
 The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity.
-In contrast, GILL, the only method in our study that does not perform latent search, struggles to maintain consistent alignment with the text instructions, resulting in more inconsistent and less coherent image sequences.
+Most methods can align images with text, however, we demonstrated that sequence decoding methods perform naturally better at aligning their output with the input sequence, and BeamDiffusion in particular achieves an optimal balance between the sequence of text instructions and the visual coherence of the generated images.
 Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process.
-Appendix A Impact of Latent Selection on Beam Search Performance
+This work shows that the latent diffusion space can be steered, offering more reliable and factual outputs.
+As future work, we plan to explore constraint decoding method that assess the likelihood of each denoise latent with respect to the target prompt.
+Solving this challenge, will improve the correlation between all elements in a sequence of scenes.
+Appendix A Implementation Details
+Base Model.
+Our method is implemented on top of Stable Diffusion v2.1 [35], a latent diffusion model for high-quality text-to-image synthesis.
+We build upon the open-source implementation provided by Stability AI, extending it to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text.
+Latent Trajectory Reuse.
+In our approach, maintaining coherence across sequential generations requires access to the intermediate latent representations produced during the generative process.
+Fortunately, the model pipeline inherently provides the full sequence of latent representations at each generation step.
+This enables us to easily extract and reuse specific latents as seed for new generations.
+Classifier Training.
+To train the contextual classifier, we employ a sequence-aware training approach that allows the model to capture dependencies between steps in the generation process.
+The classifier $\varphi$ takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and the respective step text ($B_{l-1}^{(b)}$).
+Then the classifier scores $I_{j}$ based on how well it fits within the evolving trajectory.
+During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$.
+The model is trained to minimize the cross-entropy loss:
+$\operatorname*{argmin}\sum_{d=1}^{D}\sum_{l=1}^{L}l_{d,l}\cdot\log(\varphi(I_{l},s_{l},B^{(b)}_{l-1})),$ (6)
+This contrastive training encourages the model to assign high scores to contextually appropriate continuations while penalizing incoherent or semantically unrelated ones.
+As a result, the classifier learns to distinguish subtle differences in step-to-step coherence, both visually and textually.
+Hardware.
+All experiments were ran on NVIDIA A100 GPUs with 40GB of memory.
+Appendix B Experiments Details
+B.1 Metrics and Annotations
+Automatic Metrics.
+In image sequence generation, CLIP similarity scores between visually similar frames often exceed 0.85, even for near-duplicate content.
+Wang et al.
+[49] highlight that such high CLIP-I scores can indicate semantic duplication, which poses a challenge when diversity across frames is desired.
+While CLIP-I and DINO-I are common proxies for semantic similarity, rewarding near-duplicates undermines meaningful visual progression.
+Conversely, poor alignment between images and their textual prompts is reflected in low CLIP-T scores.
+Lin et al.
+[20] address this by clipping CLIP-T values below 0.1, effectively filtering out weak text-image matches.
+Figure 4: Illustration of the limitation of CLIP-I: despite a high CLIP-I score, the result is not semantically correct or visually faithful.
+The example compares several automatic metrics (CLIP-T, CLIP-I, DINO-I, CLIP*, DINO*), highlighting discrepancies and failure cases in sequence evaluation.
+To balance these factors, we set CLIP-I scores above 0.9 and DINO-I scores above 0.85 to zero, penalizing visual redundancy.
+Similarly, we zero out CLIP-T scores below 0.1 to discourage poor alignment with the prompt.
+This clipping strategy encourages the generation of sequences that are both semantically relevant and visually diverse.
+In Figure 4, we present the automatic metric scores for a sequence across all methods.
+The results illustrate that very high CLIP-I values can lead to sequences that fail to follow the textual steps accurately, as observed in the case of GILL.
+Despite achieving strong visual similarity, the sequence diverges from the intended narrative, highlighting a key limitation of relying solely on CLIP-I .
+This example reinforces the importance of our clipping strategy: without penalizing excessively high CLIP-I values, the generation process may favor near-duplicates over meaningful visual progression, ultimately undermining semantic fidelity to the prompt.
+Table 5: Comparison of CLIP-I, DINO-I and CLIP-T across different methods with the standard deviation
+Method CLIP-I DINO-I CLIP-T
+GILL 0.55 $\pm$ 0.30 0.49 $\pm$ 0.26 0.43 $\pm$ 0.08
+StackedDiffusion 0.76 $\pm$ 0.12 0.46 $\pm$ 0.14 0.47 $\pm$ 0.09
+Image Sequence Decoding
+Greedy 0.81 $\pm$ 0.13 0.54 $\pm$ 0.14 0.48 $\pm$ 0.07
+Nucleus 0.77 $\pm$ 0.15 0.53 $\pm$ 0.15 0.48 $\pm$ 0.08
+BeamDiffusion 0.83 $\pm$ 0.14 0.62 $\pm$ 0.18 0.48 $\pm$ 0.09
+To complement average scores, we analyze the standard deviation ($\sigma$) of these metrics (Table 5).
+A low $\sigma$ reflects stable, consistent performance between diverse sequences, while a higher $\sigma$ implies variability in quality or adaptation to task complexity.
+GILL exhibits the highest $\sigma$ in CLIP-I and DINO-I, indicating its semantic similarity varies significantly between tasks, likely due to inconsistent adherence to step instructions.
+StackedDiffusion’s low $\sigma$ indicates its consistent performance across different tasks.
+BeamDiffusion strikes a balance with moderate $\sigma$ values, indicating a balance between consistency and variation across tasks.
+It is worth noting that CLIP-T generally shows lower standard deviation across models, since this metric is explicitly minimized during the diffusion model training.
+As a result, CLIP-T scores tend to be more stable regardless of the task-specific variation.
+Together, the mean and standard deviation metrics suggest BeamDiffusion offers the best combination of high semantic fidelity, text alignment, and stable adaptation across different sequences, making it the most robust method for sequential image generation.
+Human annotations.
+To facilitate the selection of the best beam search configuration, we developed a custom website specifically for human annotation.
+We found that existing tools lacked the flexibility necessary for our specific use case, especially when evaluating multiple configurations simultaneously.
+To select the best beam search configuration, 5 annotators are presented with 24 different sequences, each corresponding to a unique beam search configuration.
+In total, the 24 configurations were generated by combining three key factors:
+  • Steps Back ($m$ in Eq.
+2): This refers to how many previous steps the model uses from its latent representations to explore new candidate outputs.
+We evaluated three different settings for the number of steps back: 2, 3, and 4, which correspond to using latents from the last two, three, or four images, respectively.
+  • Latents Indexes: We tested two sets of latent configurations, 1, 3, 5, 7 and 0, 1, 2, 3, by selecting specific indexes from the latent representations generated in previous stages.
+This was done to explore how varying levels of latent detail impact the quality of the generated sequences.
+  • Beam Width ($w$ in Eq. 5): We explored four different beam widths, 2, 3, 4 and 6, which determine how many beams are retained at each step during the search process.
+Evaluating all of these at once proved to be exhausting and difficult to manage.
+To simplify the evaluation process, we adopted a elimination annotation method.
+In this approach, 5 annotators compare two configurations at a time and select the one with better overall coherence.
+If neither configuration stands out, annotators can select "both good" or "both bad" (Figure 5), in which case the computationally cheaper option advances to the next round.
+This process is repeated until we identify the best configuration.
+This method focuses on direct comparisons, making it easier to pinpoint the most effective beam search settings while also considering computational efficiency.
+Figure 5: Example of the human annotation elimination round for selecting the best beam search configuration.
+To evaluate all models performance, annotators are presented with five generated sequences side by side.
+Figure 6 illustrates an example of this comparison process, where the annotators assess the sequences to determine the most effective approach.
+Figure 6: Example of a human annotation to choose the best method.
+Gemini evaluation.
+Gemini’s selection process for annotating the best beam search configuration and the best generation method closely mirrored that of human annotators, with one key adjustment: Gemini tends to default to "both good" or "both bad" when such options are available.
+To ensure more decisive and accurate results, we removed these options from the prompt.
+We also refined our comparison methodology.
+Initially, presenting all five sequences at once led to biased evaluations, as the model disproportionately favored the last two sequences.
+To mitigate this, we adopted a pairwise evaluation strategy, comparing two sequences at a time.
+This approach resulted in more balanced and reliable assessments across all comparisons.
+Figure 7 shows the prompt used in Gemini to guide the evaluation of image sequences based on visual and semantic consistency.
+Gemini Prompt: "Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames.
+Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements.
+Semantic Consistency: Images must logically follow the described steps, clearly and accurately.
+Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic).
+Pay close attention to even minimal differences, as the sequences are very similar.
+If both sequences are equally good or bad, choose the one that is slightly better.
+Choose exactly one of the following options: - "First sequence" - "Second sequence" "
+Figure 7: Gemini prompt used to select the best beam search configuration and best method.
+B.2 Datasets
+Recipes and DIYs.
+Each manual task in both the Recipes and DIYs datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated.
+The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps.
+Most tasks feature an image for each step.
+The DIYs dataset includes a variety of tasks, from simpler projects to more complex ones, with detailed steps.
+VIST.
+We used the Descriptions of Images-in-Isolation (DII) annotations, as they allow for the generation of more precise and informative visual prompts.
+These annotations provide standalone descriptions of images, making it easier to capture their key visual elements without relying on surrounding context.
+The dataset includes 50 000 stories, offering a diverse range of visual and textual information.
+To ensure the most accurate and relevant annotation for each image, we leveraged CLIP.
+By using CLIP, we were able to select the annotation that best represented the image, improving the quality of our generated prompts and enhancing the overall alignment between text and visuals.
+Appendix C Impact of Latent Selection on Beam Search Performance
 Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search.
 By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence.
-To investigate this, we evaluate 15 tasks, each consisting of four steps, allowing us to systematically assess the effects of latent selection across a diverse set of sequences.
+This aligns with recent work on interpreting diffusion models.
+For example, Pont-Tuset et al.
+[30] ground textual descriptions in image regions, while Tang et al.
+[46] and Dewan et al.
+[5] explore how textual prompts influence generation through attention analysis and partial information decomposition.
+In contrast, our method contributes to interpretability by examining how latent space traversal, guided by beam search, affects the generation of coherent, semantically aligned image sequences.
+To systematically investigate this, we evaluate several sequences, each consisting of four steps, allowing us to assess the effects of latent selection across a diverse set of sequences.
 Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
-A.1 Impact of Latent Groups on Performance
+C.1 Impact of Latent Groups on Performance
 Annotations.
 Following the initial analysis of beam search configurations (Sec.
-5.2), we further investigated the impact of using different latent configurations on the generated sequences.
-While latents 0, 1, 2 and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance.
-For these latents, the beam search configurations yield very limited selection rates, indicating that they are rarely selected in both human and Gemini annotations.
-Specifically, for human annotations evaluated across 6 tasks, the only meaningful result is from the configuration with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%.
-Other configurations show no selection rates.
+5.3), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences.
+While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance.
+In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and Gemini annotations.
+For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%.
+All other configurations resulted in 0% selection.
+In contrast, none of the configurations showed any non-zero selection rate under Gemini annotations, reinforcing the lack of preference for these latents.
 Figure 8: Impact of using 1, 3, 5, and 7 latent configurations on image generation.
-Similarly, for Gemini annotations evaluated on the same 6 tasks, the configurations do not yield any notable results, with no selection rates.
-When the evaluation is extended to 35 tasks, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%.
+Similarly, for Gemini annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates.
+When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%.
 Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [33, 2].
 Automatic Metrics.
-Given the negative impact of higher latents on performance, as observed in human annotations, we decided to investigate the underlying reasons for this trend.
-Specifically, we explored how latents affect text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [44].
-DAAM helps interpret text-to-image diffusion models by analyzing the cross-attention maps within the denoising process.
-It generates heat maps that highlight how much each token in the prompt influences different parts of the generated image, offering a visual representation of text-image alignment.
-By applying DAAM, we aimed to determine whether higher latents distort or weaken the contribution of key tokens during generation.
-We then used DAAM to analyze the cross-attention between the input tokens and the generated image.
-DAAM enables us to track how much each remaining token influences specific regions of the image by calculating the attention patterns in the model.
-The resulting heat maps visualize this influence, with the intensity of the heat maps indicating the level of contribution each token has on different parts of the generated image.
-This approach helps us understand how well the model attends to key parts of the prompt and ensures that the most meaningful tokens are appropriately reflected in the image.
-To quantify text alignment, we calculated the average heat map maximum intensity for all remaining tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image.
+Given the observed negative impact of higher latents on performance in human annotations, we investigated how latents influence text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [46].
+DAAM interprets text-to-image diffusion models by analyzing cross-attention maps during the denoising process, generating heat maps that reveal how strongly each prompt token affects different regions of the output image.
+By applying DAAM, we examined whether higher latents distort or weaken the influence of key tokens during generation.
+The resulting heat maps allowed us to visually assess token-level contributions, helping us understand whether the model continues to attend to the most semantically important parts of the prompt under varying latent configurations..
+To quantify text alignment, we calculated the average heat map maximum intensity for all tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image.
 Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence.
@@
 The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges.
-Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523) and the 8–11 range (1.435).
+Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523).
 In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores.
@@
 While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts.
-For example, later latent groups (4-7) maintain stability in visual elements throughout the process, preserving previously established features.
+For example, later latent groups (16-19) maintain stability in visual elements throughout the process, preserving previously established features.
 However, this comes at the expense of adapting to new or changing textual instructions.
@@
 For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
-A.2 Latent Selection on Image Sequence Generation
+C.2 Latent Selection on Image Sequence Generation
 Figure 10: The average number of times that the model selected a given latent to generate the next image.
-Following our evaluation of the impact of latent groups on performance, we now examine the influence of individual latents within the groups that exhibit better text alignment (0-3).
-As shown in Figure 10, Latent 3 is used most frequently, accounting for 44.4% of selections, followed by Latent 0 at 20%.
-In contrast, Latents 1 and 2 are used less often, with Latent 1 at 11.1% and Latent 2 at 8.9%.
-This uneven distribution suggests that Latents 0 and 3 are particularly effective at capturing key features at critical points in the generation process, which beam search exploits to enhance sequence coherence.
-The ”Rand seed,” which introduces variability into the process, accounts for 15.6% of selections.
-This highlights beam search’s flexibility, allowing it to explore a range of possible outputs while preserving the overall structure of the image sequence.
-In general, these findings reinforce the importance of latent selection strategies in optimizing image generation, suggesting that models could benefit from prioritizing the most impactful latents to achieve both coherent visual sequences and stronger text-image alignment.
+Following our evaluation of the impact of latent groups on performance, we now examine the role of individual latents within the higher-performing group (Latents 0-3).
+As shown in Figure 10, Latent 3 is selected most frequently (44.4%), followed by Latent 0 (20%), while Latents 1 and 2 are used less often (11.1% and 8.9%, respectively).
+This uneven distribution suggests that Latents 0 and 3 contribute disproportionately to the generation of sequences with better text alignment.
+One possible explanation is that these latents encode more semantically relevant or visually discriminative features that align well with the text prompts.
+Their frequent selection by beam search indicates that these latents are particularly effective at resolving ambiguities or reinforcing important visual cues during sequence generation.
+The lower usage of Latents 1 and 2 may reflect their tendency to encode less salient or redundant information, which contributes less to text-conditioned coherence.
+Interestingly, the Rand seed, which injects stochasticity into the generation process, accounts for 15.6% of selections.
+This suggests that while beam search primarily exploits strong latents like 0 and 3, it also values diversity.
+The inclusion of randomized elements allows the search to escape local optima and sample from alternative yet plausible paths, enhancing the robustness of the final sequence.
 Figure 11: Heat maps illustrating cross-attention in latent diffusion across different decoding methods.
 The visualizations show variations in how each approach preserves or redistributes attention to textual cues throughout the generation process.
-Figure 11 illustrates the impact of different latent search methods on text alignment.
-The heat maps reveal how attention patterns shift based on the chosen decoding strategy, highlighting the differences in how each approach preserves or redistributes focus on textual cues throughout the image generation process.
-As shown, particularly in step 3, BeamDiffusion demonstrates superior text alignment compared to the other methods.
-Not only does it maintain a stronger connection to the input text, it also ensures better overall sequence coherence.
-This improved alignment and consistency across generations are key advantages of BeamDiffusion.
-These variations provide further insight into how latent search methods can influence both the coherence of the generated sequence and the alignment between the input text and visual output.
-Appendix B Minimizing Decoding Risk
+Figure 11 illustrates how different latent search methods affect text-image alignment.
+The heat maps show how attention patterns shift depending on the decoding strategy, revealing variations in how well each method preserves focus on key textual cues.
+Notably, in step 3, BeamDiffusion exhibits superior alignment, maintaining a stronger connection to the prompt and ensuring greater sequence coherence compared to other methods.
+While Beam Diffusion generally achieves better visual consistency and stronger text adherence in earlier steps, it slightly underperforms in step 5, failing to capture the "fireworks" mentioned in the prompt.
+This highlights how even methods with strong overall performance can exhibit localized weaknesses, underscoring the importance of analyzing alignment step by step.
+Overall, these findings highlight the significance of latent selection strategies in optimizing image generation.
+The results indicate that prioritizing the most significant latents could enhance the coherence of visual sequences and improve text alignment.
+These observations underline the influence of latent search strategies on the semantic accuracy and coherence of generated outputs.
+C.3 Effect of Cumulative Latent Access
+To investigate how access to latents from previous steps influences the quality of generation, we conducted an ablation study in which we constrain the set of available latents throughout the generation of the sequence.
+Specifically, we evaluate performance when the model is limited to latents from only the first step, then from the first two steps, and so on, up to including latents from all previous steps (i.e., full cumulative access).
+Figure 12: Visual comparison of sequences generated with cumulative latent access.
+Each row shows a sequence generated using latents from only the first step, the first two steps, and so on, up to all previous steps.
+The bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
+As shown in Figure 12, limiting the model to latents from only the first step results in weak text-image alignment and incoherent visual transitions.
+Gradually expanding the latent set to include more steps improves sequence quality, particularly when moving from two to three steps.
+However, we observe that including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
+In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case.
+This indicates that a narrow temporal window is sufficient to maintain strong alignment and visual consistency.
+By avoiding the computational cost of considering all previous steps, BeamDiffusion finds a favorable trade-off between quality and efficiency.
+Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary.
+Strategies such as BeamDiffusion, which take advantage of a limited but informative temporal context, can produce high-quality sequences while significantly reducing computational overhead.
+These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents without the need for full sequence access.
+Appendix D Minimizing Decoding Risk
 The generation process using diffusion models begins with a seed that determines the initial conditions of the model.
@@
 The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$ may have one scene in the middle of the sequence that should be independent.
-To address this, for steps $j>1$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
-Appendix C Annotation Tasks
-Human annotations.
-To facilitate the selection of the best beam search configuration, we developed a custom website specifically for human annotation.
-We found that existing tools lacked the flexibility necessary for our specific use case, especially when evaluating multiple configurations simultaneously.
-To select the best beam search configuration, 5 annotators are presented with 24 different sequences, each corresponding to a unique beam search configuration.
-In total, the 24 configurations were generated by combining three key factors:
-  • Steps Back: This refers to how many previous steps the model uses from its latent representations to explore new candidate outputs.
-We evaluated three different settings for the number of steps back: 2, 3, and 4, which correspond to using latents from the last two, three, or four images, respectively.
-  • Number of Latents: We tested two sets of latent configurations {1, 3, 5, 7} and {0, 1, 2, 3} to explore how varying levels of latent detail impact the quality of the generated sequences.
-  • Beam Width: We explored four different beam widths {2, 3, 4, 6} which determine how many candidate sequences are retained at each step during the search process.
-Evaluating all of these at once proved to be exhausting and difficult to manage.
-To simplify the evaluation process, we adopted a ”battle-style” annotation method.
-In this approach, 5 annotators compare two configurations at a time and select the one they find has better overall coherence.
-If neither configuration stands out, annotators can select ”both good” or ”both bad” (Figure 12), in which case the computationally cheaper option advances to the next round.
-This process is repeated until we identify the best configuration.
-This method focuses on direct comparisons, making it easier to pinpoint the most effective beam search settings while also considering computational efficiency.
-Figure 12: Example of the human annotation battle round for selecting the best beam search configuration.
-To evaluate all models, 10 annotators are presented with four generated sequences side by side: one from the top-performing beam search configuration (selected during the initial annotations) and three others from alternative decoding methods, such as greedy search, nucleus sampling, and GILL.
-Figure 13 illustrates an example of this comparison process, where the annotators assess the sequences to determine the most effective approach.
-Figure 13: Example of a human annotation to choose the best decoding method.
-Gemini evaluation.
-Figure 14 shows the prompt used in Gemini to annotate the best configuration and the best method.
-This prompt was used to guide the evaluation of image sequences based on visual and semantic consistency.
-Gemini Prompt: ”Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames.
-Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements.
-Semantic Consistency: Images must logically follow the described steps, clearly and accurately.
-Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic).
-Pay close attention to even minimal differences, as the sequences are very similar.
-If both sequences are equally good or bad, choose the one that is slightly better.
-Choose exactly one of the following options: - ”First sequence” - ”Second sequence” ”
-Figure 14: Gemini prompt used to select the best beam search configuration and best method.
-Appendix D Datasets
-Recipes and DIYs.
-Each manual task in both the Recipes and DIYs datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated.
-The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps.
-Most tasks feature an image for each step.
-The DIYs dataset includes a variety of tasks, from simpler projects to more complex ones, with detailed steps.
-VIST.
-We used the Descriptions of Images-in-Isolation (DII) annotations, as they allow for the generation of more precise and informative visual prompts.
-These annotations provide standalone descriptions of images, making it easier to capture their key visual elements without relying on surrounding context.
-The dataset includes 50 000 stories, offering a diverse range of visual and textual information.
-To ensure the most accurate and relevant annotation for each image, we leveraged CLIP.
-By using CLIP, we were able to select the annotation that best represented the image, improving the quality of our generated prompts and enhancing the overall alignment between text and visuals.
-Appendix E Challenges
+To address this, for steps $s_{j>1}$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
+Appendix E Broader Impacts and Challenges
 In the development and evaluation of the BeamDiffusion method, we encountered several challenges that required careful consideration and innovation.
@@
 In this section, we outline the key challenges we faced and the strategies we employed, while acknowledging that not all of them have been fully addressed.
-Captions.
-One of the challenges encountered in this process is generating captions that accurately describe the steps involved in the sequence.
+E.1 Impacts
+This work proposes a method to improve generative pretraining by incorporating subject-aware mutual information maximization, aiming to produce more contextually aligned and coherent sequences of images.
+The ability to generate visually consistent image sequences has positive applications in areas such as animation, visual storytelling, educational media, and accessibility tools (e.g., for generating illustrative content from text).
+However, as with many generative models, there are potential negative societal impacts.
+Enhanced coherence and subject fidelity could be misused to create misleading or deceptive visual content, such as more convincing visual disinformation, propaganda, or synthetic media (e.g., deepfakes).
+E.2 Contextual Scene Descriptions
+One of the challenges encountered concerns the generation of captions that accurately describe the steps involved in the sequence.
 Often, a single step may consist of multiple sub-steps, each contributing to the overall generation.
 This makes it difficult to write a single concise caption that captures the full scope of each step.
-The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each stage but also conveys the progression and complexity of the entire sequence.
+The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each step but also conveys the progression and complexity of the entire sequence.
 Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
-Domains.
+To address this challenge, we leverage Gemini [34], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context.
+Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
+$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (7)
+Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps.
+E.3 Domains.
 Identifying a domain that effectively highlights the advantages of the BeamDiffusion method, particularly the benefits of beam search, is still a challenge.
@@
 Finding a domain that not only aligns with the method’s capabilities but also provides a clear and compelling demonstration of its improvements remains an ongoing challenge.
-Complexity of the generation.
+E.4 Complexity of the generation.
 The BeamDiffusion model generates multiple candidate images at each step by exploring a range of latents.
-We analyze the complexity of the model for a steps back of $n$ steps and $k$ latents explored.
+We analyze the complexity of the model for a steps back of $m$ steps and $N$ latents explored.
 In the initial steps, the complexity is high because pruning begins only after the second step.
-For each beam, $k\times n$ latents are explored, and the search space grows exponentially, resulting in significant computational cost in the early stages.
-During the first two steps, the complexity grows with the number of candidates and latents, proportional to $O((k\times n)^{j})$ for $j\leq 2$.
+For each beam, $N\cdot m$ latents are explored, and the search space grows exponentially, resulting in significant computational cost in the early stages.
+During the first two steps, the complexity grows with the number of candidates and latents, proportional to $\mathcal{O}(N\cdot m)^{j})$ for $j\leq 2$.
 Specifically, for $j=1$, the complexity is proportional to the number of explored random seeds.
-After the second step, pruning reduces the number of candidates to the most promising $w$ beams, decreasing the complexity to $O(k\times n\times w)$.
-To further reduce complexity, a heuristic based on text relevance can be used to predict the top $T$ most relevant past images.
-By restricting the search space to the latents of these top $T$ candidates, the model improves efficiency by prioritizing contextually relevant images, particularly in the early steps.
+After the second step, pruning reduces the number of candidates to the most promising $w$ beams, decreasing the complexity to $\mathcal{O}(N\cdot m\cdot w)$.
+To further reduce complexity, a heuristic based on text relevance can be used to predict the top $P$ most relevant past images.
+By restricting the search space to the latents of these top $P$ candidates, the model improves efficiency by prioritizing contextually relevant images, particularly in the early steps.
 While the initial three steps remain computationally demanding due to the expansive search space, pruning combined with this heuristic can significantly lower the computational cost by refining the candidate set.
 Appendix F Application Examples
-Images and other visual aids are essential tools that significantly improve comprehension in various fields.
-Whether while breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that makes the material easier to understand [26].
-From technical procedures and instructional design to storytelling and educational content, the use of multiple-scene images allows the reader to grasp information more deeply and intuitively, transforming even the most complex or dense topics into something more accessible [39, 11].
-However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence is a major challenge  [25].
+This section presents additional visual examples and qualitative results that complement the main paper.
+Images and other visual aids are essential tools that significantly improve comprehension across a variety of domains.
+Whether breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that enhances understanding [24].
+From technical procedures and instructional design to storytelling and educational content, multi-scene image sequences help readers grasp information more intuitively, transforming even the most complex or dense topics into something more accessible.
+However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence remains a major challenge [23].
+In this appendix, we include expanded examples to further illustrate how our model addresses this issue.
+These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes, visual storytelling, and DIY tasks.
+Together, these examples highlight the strengths and limitations of current approaches in achieving textual alignment and visual consistency in sequential image generation.
 Figure 13: Visualization of cross-attention in latent diffusion across different sets of latents.
-The heat maps reveal how attention patterns evolve at various stages, highlighting the alignment between text prompts and distinct latent representations.
+The heat maps reveal how attention patterns evolve at various sequence steps, highlighting the alignment between text prompts and distinct latent representations.
 F.1 Beam Search Tree
@@
 Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4.
-This demonstrates how the latent search process incorporates context from various stages, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
+This demonstrates how the latent search process incorporates context from various steps, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
 F.2 Sequence Examples
-To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 20-28.
+To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 18-26.
 These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
-In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 20, 20, 24, and 24.
+In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 18, 18, 22, and 22.
 They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps.
 This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt.
-In contrast, GILL exhibits a different behavior.
-While it maintains high visual consistency, with all images in the sequence being visually similar, it fails to accurately follow the step-by-step textual instructions.
-GILL tends to generate images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete.
+In contrast, GILL and StackedDiffusion exhibit a different behavior.
+While both maintain high visual consistency, generating sequences in which the images are visually similar, they struggle to accurately follow the step-by-step textual instructions.
+GILL, in particular, tends to produce images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete.
+StackedDiffusion performs slightly better than GILL in this regard, capturing some of the step-wise changes, but still often fails to fully reflect the evolving textual input across the sequence.
 Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency.
-As shown in Figures 20, 20, 24, and 24, it generally produces the most coherent sequences compared to other methods.
+As shown in Figures 18, 18, 22, and 22, it generally produces the most coherent sequences compared to other methods.
 In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps.
@@
 While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches.
-However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 28, 28, and 28.
+However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 26, 26, and 26.
 The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence.
 While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
-Figure 17: Example of the Recipes domain task ”Vegan Basil Pesto and Sun-Dried Tomato Pinwheels”, with a sequence of 4 steps.
-Figure 18: Example of the Recipes domain task ”Tangy lemon sheet cake”, with a sequence of 4 steps.
-Figure 19: Example of the Recipes domain task ”Zucchini, Pecorino & Mint Salad”, with a sequence of 4 steps.
-Figure 20: Example of the Recipes domain task ”Bone Broth”, with a sequence of 3 steps.
-Figure 21: Example of the out-of-domain VIST story ”Klamath Project Construction”, with a sequence of 5 steps.
-Figure 22: Example of the out-of-domain VIST story ”Yosemite Weekend”, with a sequence of 5 steps.
-Figure 23: Example of the out-of-domain VIST story ”Fireworks - 4th July in Washington, DC”, with a sequence of 5 steps.
-Figure 24: Example of the out-of-domain VIST story ”Independence Day Parade”, with a sequence of 5 steps.
-Figure 25: Example of the out-of-domain DIY task ”How to Clean Silk Rugs”, with a sequence of 2 steps.
-Figure 26: Example of the out-of-domain DIY task ”How to Polish Wood”, with a sequence of 3 steps.
-Figure 27: Example of the out-of-domain DIY task ”How to Preserve Basil”, with a sequence of 6 steps.
-Figure 28: Example of the out-of-domain DIY task ”How to Clean Natural Stone”, with a sequence of 5 steps.
+Figure 15: Example of the Recipes domain task "Vegan Basil Pesto and Sun-Dried Tomato Pinwheels", with a sequence of 4 steps.
+Figure 16: Example of the Recipes domain task "Tangy lemon sheet cake", with a sequence of 4 steps.
+Figure 17: Example of the Recipes domain task "Zucchini, Pecorino & Mint Salad", with a sequence of 4 steps.
+Figure 18: Example of the Recipes domain task "Bone Broth", with a sequence of 3 steps.
+Figure 19: Example of the out-of-domain VIST story "Klamath Project Construction", with a sequence of 5 steps.
+Figure 20: Example of the out-of-domain VIST story "Yosemite Weekend", with a sequence of 5 steps.
+Figure 21: Example of the out-of-domain VIST story "Fireworks - 4th July in Washington, DC", with a sequence of 5 steps.
+Figure 22: Example of the out-of-domain VIST story "Independence Day Parade", with a sequence of 5 steps.
+Figure 23: Example of the out-of-domain DIY task "How to Clean Silk Rugs", with a sequence of 2 steps.
+Figure 24: Example of the out-of-domain DIY task "How to Polish Wood", with a sequence of 3 steps.
+Figure 25: Example of the out-of-domain DIY task "How to Preserve Basil", with a sequence of 6 steps.
+Figure 26: Example of the out-of-domain DIY task "How to Clean Natural Stone", with a sequence of 5 steps.

### v3

@@
-Latent Beam Diffusion Models for Decoding Image Sequences
+Latent Beam Diffusion Models for Generating Visual Sequences
 Abstract
-While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency in image sequences.
-Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent frames.
+While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency when generating image sequences.
+Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent images.
 We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding.
-Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions.
+In contrast to earlier methods that rely on fixed latent priors, our method dynamically samples past latents to search for an optimal sequence of latent representations, ensuring coherent visual transitions.
 As the latent denoising space is explored, the beam search graph is pruned with a cross-attention mechanism that efficiently scores search paths, prioritizing alignment with both textual prompts and visual context.
 Human and automatic evaluations confirm that BeamDiffusion outperforms other baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment.
-All data and code implementations will be released after publication.
+[^1] Data and Code: huggingface.co/Gui28F/BeamDiffusion
 1 Introduction
@@
 This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence.
-This dynamic can result in sequences that lack coherence and visual consistency over time [2].
-We propose to utilize beam search as a potential solution to this problem.
-Beam search is a well-established technique in text generation tasks like machine translation [14, 51, 17, 37], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions.
+Figure 1: BeamDiffusion leverages search in the latent denoising space to generate a coherent sequence of images from stepwise text instructions.
+The model conditions the generation of each new image $I_{j}$ on latents from previous steps $s_{j}$, forming a chain that carries forward information from past generations.
+At each step, multiple candidate latents are explored, the worst candidates are discarded, and the best ones are kept for the next step.
+We propose to research beam search as a scientifically well-grounded solution to this problem.
+Beam search is a core text generation technique for tasks like machine translation and standard inference [16, 54, 19, 41], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions.
 However, despite its success in other domains, it has been barely explored in this context, leaving its potential untapped.
-In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues.
-In this work, we explore beam search for latent prior selection in image sequence generation.
-While previous research has used latent representations as priors to improve sequence consistency [2, 33], it remains unclear which specific representations yield the best overall sequence.
-We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences.
-As illustrated in Figure 1(a), our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
-As the beam diffusion algorithm traverses the latent space, the number of hypothesis grow and the least probably search paths are pruned to keep a fixed number of candidate paths.
-To this end, we incorporate a cross-attention mechanism [19] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps.
-This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
-We evaluated the effectiveness of BeamDiffusion in three domains: the Recipes domain and in the out-of-domain DIY and Visual Storytelling.
+In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues, as illustrated in Figure 1.
+In this work, we revisit the latent diffusion framework with a focus on structured sequence generation.
+We aim to dynamizally sample the latent space of past diffusion processes and propose the BeamDiffusion method for improving multi-image visual consistency by leveraging shared latent representations in a principled way.
+BeamDiffusion explores the latent denoising trajectories with a beam search strategy for ensuring a conditional dependency between diffusion processes.
+As illustrated in Figure 2(a), our approach leverages a sampling strategy in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images.
+As the beam diffusion algorithm traverses the latent space, the number of hypothesis grows and the search paths with the lowest probability are pruned to keep a fixed number of candidate paths.
+To this end, we incorporate a cross-attention mechanism [21] that evaluates and scores candidate image paths based on their alignment, providing a reliable quality measure for each path, allowing us to retain the best ones while pruning the least likely.
+To test the generalization and robustness of BeamDiffusion, we evaluated it in two diffusion architectures (LDM [39] and DiT [2]) and two domains (recipes and visual storytelling domains).
 Results obtained with human annotations, automatic evaluations and LLM-as-a-judge, all showed that the proposed approach outperforms baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
@@
 Given a sequence of steps $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each sequence step $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps of the sequence.
-While Latent Diffusion Models (LDMs) [35] can generate high-quality images in a compressed latent space, this latent space is designed to generate images independently, starting from a randomly initialized latent $z_{T}$.
-LDMs encode an input image $I$ into thelatent space representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [35].
-The reverse diffusion process is modeled by a U-Net backbone [36], $\epsilon_{\theta}$, which incorporates a cross-attention mechanism [47] to condition the denoising iterations on an external input, e.g.
-the output embeddings of a text encoder $\tau_{\theta}(s)$.
-LDMs are trained to iteratively estimate a denoised variant of $z$, with the loss
-$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
-This loss evidences the lack of awareness of other denoising processes, making it challenging to ensure consistency acrosscorrelated prompts or previous denoising processes.
-This is particularly challenging in non-linear narratives [42], where maintaining character identity, lighting, and environmental details are crucial.
-To address these challenges, prior work [32, 21, 45, 2, 33] reuse latent denoising iterations from earlier steps in the sequence, paving the way to a key question: how to determine the optimal set of denoising latents that maximizes coherence across different generations?
-The answer lies in a combinatorial search space, that BeamDiffusion can traverse with a search strategy to determine an optimal solution, as we will discuss next.
+Diffusion Models generate high-quality images in a compressed latent space, but typically treat each image independently from a randomly initialized latent $z_{T}$.
+Within this framework, Latent Diffusion Models (LDMs) [39] and Diffusion Transformers (DiTs) [32] share a common structure: both encode an input image $I$ into a latent representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [39].
+The reverse diffusion process is performed by a denoising backbone $\epsilon_{\theta}$, which can be a U-Net [40] (for LDMs) or a transformer [32] (for DiT), and incorporates cross-attention to condition the denoising iterations on an external input, e.g., the output embeddings of a text encoder $\tau_{\theta}(s)$.
+Both models are trained to iteratively estimate a denoised latent with the loss:
+$L_{\text{diff}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
+This formulation highlights that each denoising process is agnostic to other generations, making it challenging to maintain coherence across a sequence of related prompts.
+This is particularly important in non-linear narratives [46], where preserving character identity, lighting, and environmental details is crucial.
+To address these challenges, prior work [36, 23, 49, 3, 37] has proposed reusing latent denoising iterations from earlier steps, raising the key question: how can we determine the optimal set of latent states that maximizes coherence across different generations?
+In the following subsections, we introduce BeamDiffusion (Sec. 3.1), detail latent space exploration with diffusion beams (Sec. 3.2), and outline our beam pruning strategy (Sec. 3.3).
 3.1 Beam Diffusion Model
-To address the limitations of standard LDMs in generating coherent image sequences, we introduce the BeamDiffusion model.
+To address the limitations of standard diffusion models in generating coherent image sequences, we introduce the BeamDiffusion model.
 The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space.
-BeamDiffusion incorporates latent representations from previously generated images.
+BeamDiffusion incorporates latent representations from previously generated images, Figure 2(a).
 These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence.
 This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the next scene description.
-In the remainder of this article, the term step will always refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
-3.1.1 Latent Space Exploration with Diffusion Beams
+In this article, the term step will refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
+3.2 Latent Space Exploration with Diffusion Beams
 To generate coherent and diverse image sequences, we explore the latent space of diffusion models using a beam search-inspired strategy.
 For each sequence step $j$ we generate a set of candidate images $\mathcal{I}_{j}=\{I_{j}^{(1)},I_{j}^{(2)},\dots\}$, where each $I_{j}^{(k)}$ is conditioned on a different beam path.
-Algorithm 1 BeamDiffusion Algorithm
-1:  Input parameters: $\{s_{1},\ldots,s_{N}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
-2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$triangleright Beam initialization with $r$ different images.
-3:  for $j=2$ to $L$ do
-4:     for all beam $b\in\hat{B}_{j-1}$ do
-5:        $\mathcal{Z}_{j}\leftarrow$$\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$triangleright Read denoising latents from previous steps.
-6:        for all latent $z\in\mathcal{Z}_{j}$ do
-7:           $I_{j}\leftarrow\text{SD}(s_{j},z)$triangleright Generate candidate a image.
-8:           $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$triangleright Extend beam with candidate image.
-9:        end for
-10:     end for
-11:     $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$triangleright Score the new image of the beam.
-12:     $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$triangleright Prune beams and keep top $w$ beams.
-13:  end for
-14:  return  best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
 Sampling the Denoising Latent Space.
@@
 This step establishes a dependency between generations, where latents from earlier steps guide the synthesis of visually coherent transitions over time, as illustrated in Figure 2.
+Algorithm 1 BeamDiffusion Algorithm
+1:  Input parameters: $\{s_{1},\ldots,s_{L}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
+2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$ $\triangleright$ Beam initialization with $r$ different images.
+3:  for $j=2$ to $L$ do
+4:   for all beam $b\in\hat{B}_{j-1}$ do
+5:    $\mathcal{Z}_{j}\leftarrow$ $\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$ $\triangleright$ Read denoising latents from previous steps.
+6:    for all latent $z\in\mathcal{Z}_{j}$ do
+7:     $I_{j}\leftarrow\text{SD}(s_{j},z)$ $\triangleright$ Generate candidate a image.
+8:     $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$ $\triangleright$ Extend beam with candidate image.
+9:    end for
+10:   end for
+11:   $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$ $\triangleright$ Score the new image of the beam.
+12:   $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$ $\triangleright$ Prune beams and keep top $w$ beams.
+13:  end for
+14:  return best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
 Diffusion Beams.
 Inspired by beam search [53], which maintains and extends the most promising candidates at each sequence step, we introduce diffusion beams to structure sequential image generation.
-At each step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step $j-1$ and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$.
+At each sequence step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step j1j-1 and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$.
 The resulting beams are defined as
@@
 1, line 8, enabling the model to explore diverse yet coherent visual narratives over time.
-3.1.2 Pruning Diffusion Beam Paths
+3.3 Pruning Diffusion Beam Paths
 Tracking multiple diffusion beams enables BeamDiffusion to maintain visual coherence while exploring diverse continuations across a sequence.
@@
 Training.
-The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well align an image is with the previously sequence of steps and images.
+The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well aligned an image is with the previously sequence of steps and images.
 For a given sequence $t$ and sequence step $j$, the ground-truth label $l_{t,j}$ indicates whether image $I_{j}$ belongs to the sequence given its prompt $s_{j}$ and context $B_{j-1}^{(b)}$.
@@
 The model minimizes the cross-entropy loss encouraging the model to promote candidates that best match the sequence’s semantic and visual dynamics.
-See appendix A for implementation details.
+The classifier is trained on the Recipes [3] and VIST [15] datasets.
+See appendix H for implementation details.
 4 Experimental Setup
-In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences.
-We outline the datasets used, identify other baseline methods, and explain the human, automatic and LLM-as-a-judge (Gemini [34]) evaluations focused on semantic and visual consistency across generated sequences.
+In this section, we describe the experimental setup for evaluating BeamDiffusion’s performance in generating coherent image sequences.
+We first outline the different decoding strategies that are considered.
+We then present the datasets and baseline methods used, and describe the evaluation protocols, which include human judgments, automatic metrics, and LLM-as-a-judge evaluations using Gemini 2.5 [6], all designed to assess semantic and visual consistency across generated sequences.
+Decoding strategies.
+Our proposed method, BeamDiffusion, uses a beam search over latent sequences to generate coherent image sequences from textual prompts.
+To evaluate the effectiveness of this decoding strategy, we also compare it against two alternative approaches applied to the same model: Greedy / CoSeDlen=1 [37], which selects at each step the top-scoring image according to CLIP similarity from the first few latents of the previous step, focusing on immediate best-match selection; and Nucleus Sampling [13], a probabilistic strategy that samples images from the subset whose cumulative CLIP probability exceeds a threshold $p$, encouraging diversity while still prioritizing high-probability outcomes.
+All strategies are applied to the same sequences of prompts using two underlying diffusion backbones, FLUX.1-dev [2] and SD2.1 [39], allowing a direct comparison of their ability to maintain sequence-level coherence across different models.
 Datasets.
-We used a dataset consisting of publicly available manual tasks in the Recipes and DIY domains [2], as well as the Visual Storytelling (VIST) dataset [13].
-Recipes serve as the in-domain dataset, while DIY tasks and VIST are out-of-domain.
-These out-of-domain datasets allow us to assess the generalizability of our model across different task domains.
-More details can be found in Appendix B.2.
+We used publicly available manual tasks in the Recipes domain [3], along with a small portion of the Visual Storytelling (VIST) dataset [15], using an equal number of tasks from each.
+Recipes provide detailed, step-by-step instructions, while the sampled VIST tasks offer additional diversity for training the classifier.
+This combination allows us to evaluate BeamDiffusion’s performance on both familiar and novel task types.
+Together, these datasets enable a thorough assessment of sequence coherence, semantic alignment, and visual consistency.
+More details can be found in Appendix F.
 Baselines.
-To evaluate the performance of our image sequence generation method, we compared BeamDiffusion with SD2.1 [35] against two other competing approaches to generate image sequences: GILL [18] and StackedDiffusion [26].
-Additionally, we considered two decoding strategies: Greedy and Nucleus Sampling [11].
-As these decoding strategies are novel in the context of image generation, we provide a brief explanation of each.
-All methods used 50 denoising iterations.
-In the Greedy approach, a new image is generated at each step using the first five latents from the previous image, with the image that scores highest overall being selected.
-Although this is similar to CoSeDlen=1 [33], it differs in that it relies exclusively on CLIP for selection without any learned module, whereas CoSeD incorporates a trained selection mechanism.
-In contrast, Nucleus Sampling introduces diversity by sampling from the subset of images whose cumulative CLIP-assigned probability exceeds a predefined threshold $p$.
-Rather than always selecting the most likely image, this method favors high-probability outcomes while allowing for variation in the generated sequence.
+To assess the performance of BeamDiffusion, we compared against several competitive baselines that aim to generate coherent image sequences from text: GILL [20], StackedDiffusion [29], StoryDiffusion [55], and One-Prompt-One-Story (1P1S) [24].
 Human Evaluation.
@@
 2).
-The winning configuration is compared against other configurations until a new configuration is selected as better.
-The selection criteria focused on semantic alignment and visual consistency.
-Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
+The winning configuration is selected iteratively by comparing different configurations and preserving always the better one.
+See Appendix B for the detailed results.
 Having chosen the optimal configuration, we compared BeamDiffusion to other methods.
@@
 Annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency with the possibility of choosing multiple options or none.
-For further information, see Appendix B.1.
-Gemini Evaluation.
-To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines.
+For further information, see Appendix E.
+LLM-as-a-judge Evaluation.
+To further validate our findings and annotate a wider range of tasks, we adopt the LLM-as-a-judge paradigm, leveraging Gemini, to evaluate both configurations and baselines.
 Its selection process closely mirrored that of human annotators.
-For details, please refer to Appendix B.1.
-Human-Gemini Agreement.
-We conducted an alignment assessment to measure the agreement between human annotators and Gemini.
-For this, human annotators were asked to rate 16 Recipes, 10 DIYs, and 10 VIST tasks.
-These ratings were then compared to the evaluations made by Gemini.
+For details, please refer to Appendix E.
+Human and LLM-as-a-judge Agreement.
+We conducted an alignment assessment to measure agreement between human annotators and the LLM-as-a-judge.
+Seven annotators rated 16 Recipes and 10 VIST tasks, which we compared to the model’s evaluations.
 To quantify the agreement between the two, we computed Fleiss’ kappa [8], obtaining a value of 70%, which indicates a substantial level of agreement.
-Based on this strong alignment, we extended the evaluation to the remaining sequences using Gemini.
+Building on this strong alignment, we employed the LLM-as-a-judge to evaluate the remaining sequences.
 Automatic Metrics.
-Following prior work [2, 33, 21], we evaluate each method using: (1) Visual Consistency, based on average image similarity via DINO [27] (DINO-I) and CLIP [31] (CLIP-I); and (2) Semantic Alignment, using CLIP (CLIP-T) to measure similarity between the current image and the average of all step texts up to and including the current one.
-To place the scores on a comparable scale, we normalize CLIP-I, CLIP-T and DINO by their maximum.
-Finally, we compute CLIP* as the product between the CLIP-I and CLIP-T, and DINO* as the product between the DINO-I and CLIP-T, this formulations balances visual and semantic quality, while emphasizing semantic alignment, which is critical for accurately reflecting the stepwise input text.
-See Appendix B.1 for details.
+Following prior work [3, 37, 23], we evaluate each method using two complementary criteria: (1) Visual Consistency, measured via average image similarity using DINO [30] (DINO-I) and CLIP [35] (CLIP-I); and (2) Semantic Alignment, computed with CLIP (CLIP-T) by comparing each image to the average of all step texts up to and including that step.
+To ensure comparability across metrics, we normalize CLIP-I, CLIP-T, and DINO-I by their respective maxima.
+We further define CLIP* as the product of CLIP-I and CLIP-T, and DINO* as the product of DINO-I and CLIP-T, balancing visual and semantic quality while emphasizing alignment with the stepwise input text.
+To better capture real performance, the results for DINO-I, CLIP-I, CLIP-T, CLIP*, and DINO* are reported as win percentages rather than mean values.
+This avoids overestimating methods that perform very well on some sequences but poorly on others, which could otherwise produce artificially high averages despite inconsistent quality.
+In addition, we adopt metrics from StackedDiffusion [29] to evaluate task-specific aspects of the generated sequences: Goal Faithfulness, measuring alignment of the final image with the overall textual goal; Step Faithfulness, measuring alignment of each intermediate image with its corresponding step text; and Cross-Image Consistency, capturing visual coherence across images in the sequence.
+These metrics complement the previous ones by explicitly assessing semantic fidelity and sequence-level consistency.
+See Appendix E for further details.
 5 Results and Discussion
-In this section, we describe the performance of BeamDiffusion based on both human and automatic evaluations.
-We begin by analyzing the human evaluation, proceed to discussing the automatic evaluations, then focus on specifics of our approach such as hyperparameters and non-linear decoding characteristics, and finally provide a qualitative analysis.
-Additional statistical information about the results can be found in Appendix B.1.
+In this section, we evaluate the effectiveness of BeamDiffusion through three complementary approaches: human evaluations, LLM-as-a-judge, and automatic metrics.
+We begin by comparing decoding strategies, selecting the method that achieves the highest human preference scores for subsequent experiments.
+We then highlight key performance aspects of BeamDiffusion, including visual and semantic consistency, and conclude with a qualitative analysis.
+Additional details on the annotation setup are provided in Appendix E.
+5.1 Comparison of Decoding Methods
+To evaluate the effectiveness of different decoding strategies, we compare our proposed method, BeamDiffusion, against Greedy/CoSeDlen=1 [37] and Nucleus Sampling [13] using the same sequences of prompts with two types of diffusion backbones.
+Table 1: Comparison of human (H) and LLM-as-a-judge (L) evaluations of decoding strategies across two backbones.
+FLUX.1-dev SD2.1
+Method H (%) L (%) H (%) L (%)
+Greedy/CoSeDlen=1 [37] 16.35 10.00 11.11 10.00
+Nucleus Sampling [13] 24.04 16.67 33.33 26.67
+BeamDiffusion 59.61 73.33 55.56 63.33
+As shown in Table 1, BeamDiffusion consistently achieves the highest preference scores according to both human annotators (H) and LLM-as-a-judge (L).
+In particular, BeamDiffusion substantially outperforms Greedy decoding and Nucleus Sampling, indicating that beam-based decoding produces sequences that are not only more coherent but also better aligned with human and automatic judgments.
+While Nucleus Sampling improves over Greedy decoding by encouraging diversity, its performance remains notably below that of BeamDiffusion in both evaluation settings.
+The agreement between H and L further strengthens the evidence for the robustness of our decoding strategy.
+Given these results, we adopt BeamDiffusion as our default decoding method in subsequent experiments (see Appendix C.4 for qualitative and automatic evaluation results and Appendix B for hyperparameter configurations).
+Table 1 also compares two diffusion architectures: Stable Diffusion (SD2.1) and Diffusion in Transformers (FLUX-1-dev).
+BeamDuffusion exhibits consistent performance with both architecture, with Diffusion in Transformers outperforming SD2.1.
+Hence, in the following experiments we will use the Diffusion in Transformers architecture.
+Table 2: Frequency of method selection by human annotators (H) and the LLM-as-a-judge (L).
+Recipes VIST
+Method H (%) L (%) H (%) L (%)
+GILL [20] 0.00 0.00 2.74 0.00
+StackedDiffusion [29] 0.00 3.33 0.00 0.00
+StoryDiffusion [55] 22.58 20.00 36.99 41.38
+1P1S [24] 12.90 3.33 19.18 10.34
+BeamDiffusion 64.52 73.33 41.10 48.28
 5.2 General Results
-Table 1 compares all methods across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain).
-BeamDiffusion consistently outperforms other methods.
-In the context of Recipes, BeamDiffusion is selected 50.0% of the time by humans and 37.6% by the LLM, with Nucleus and Greedy following.
-In out-of-domain areas, it is chosen once more 50.0% of the time by humans and 30.0% by the LLM-judge for the DIY domain, and 58.3% for humans and 34.9% by the LLM for the VIST dataset.
-Among sequence generation methods, GILL consistently underperforms across all domains, receiving at most 3.3% of human votes and 16.1% from the LLM in VIST.
-StackedDiffusion performs better than both Greedy and GILL, but still lags behind BeamDiffusion, with 20.0% of human selections and 9.8% of LLM-judge selections in the Recipes domain.
-Overall, BeamDiffusion excels in Recipes and VIST, however performance in DIY remains challenging, with smaller gaps between methods due to the domain’s inherent complexity.
-Table 1: Frequency of method selection by human annotators(H) and Gemini(G).
-Recipes DIY VIST
-Method H (%) G (%) H (%) G (%) H (%) G (%)
-GILL [18] 0.0 2.0 0.0 20.0 3.3 16.1
-StackedDiffusion [26] 20.0 9.8 12.5 15.6 0.00 3.2
-Image Sequence Decoding
-Greedy/CoSeDlen=1 [33] 0 27.6 0 23.3 0 26.3
-Nucleus Sampling [11] 30.0 23.0 37.5 27.7 8.3 26.6
-BeamDiffusion 50.0 37.6 50.0 30.0 58.3 34.9
+Table 2 compares the sequence generation methods on both the Recipes and VIST datasets.
+BeamDiffusion dominates in both domains, leading Recipes with 64.52% (H) and 73.33% (L), and VIST with 41.10% (H) and 48.28% (L).
+StoryDiffusion consistently ranks second, with 22.58% (H) and 20.00% (L) in Recipes, and 36.99% (H) and 41.38% (L) in VIST.
+In contrast, 1P1S receives moderate preference (12.90% H in Recipes, 19.18% H in VIST), while GILL and StackedDiffusion record the lowest selection rates.
+These results show that BeamDiffusion achieves the strongest performance across settings, with especially large margins in Recipes where BeamDiffusion surpasses the next-best method by over 40% for human preference.
+The consistently low selection rates of GILL and StackedDiffusion highlight that these methods are rarely preferred by either humans or LLM-as-a-judge, in contrast to the stronger performance of BeamDiffusion and StoryDiffusion.
+Overall, the findings highlight that effective search and selection strategies are central to producing coherent and engaging narratives, enabling models to maintain quality even in complex domains such as VIST.
+Table 3: Comparison of different methods across embedding-based sequence quality metrics.
+CLIP-I, DINO-I, CLIP-T, CLIP*, and DINO* are reported as win percentages, while Goal Faithfulness, Step Faithfulness, and Cross-Image Consistency are reported as absolute scores.
+[…]
+GILL [20] 0.00 0.00 0.00 0.00 0.00 0.67 0.52 0.54
+StackedDiffusion [29] 25.00 25.00 5.00 25.00 25.00 0.60 0.67 0.57
+StoryDiffusion [55] 20.00 25.00 35.00 25.00 20.00 0.78 0.87 0.61
+1P1S [24] 25.00 30.00 10.00 10.00 25.00 0.75 0.83 0.57
+BeamDiffusion 30.00 20.00 50.00 40.00 30.00 0.82 0.92 0.61
 5.3 Automatic Metrics
-Table 2 shows that BeamDiffusion achieves the highest CLIP* and DINO* scores (0.40 and and 0.30), demonstrating a strong balance between image consistency and semantic alignment.
-BeamDiffusion excels in maintaining a good compromise between text grounding and image similarity achieving the best result in image consistency (CLIP-I of 0.83 and DINO-I of 0.62).
-Similarly, for CLIP-T, BeamDiffusion achieves the highest score for image-text alignment (CLIP-T of 0.48).
-However, in this case, the highest result is shared, as Greedy and Nucleus obtain the same score.
-This outcome is expected because Greedy and Nucleus utilize CLIP-T to select their outputs by aligning the generated sequence with the text.
-In terms of sequence generation methods, StackedDiffusion performs better than GILL in all metrics but still lags behind BeamDiffusion and the other decoding methods.
-Table 2: Comparison of CLIP-I, DINO-I and CLIP-T across different methods
-Method CLIP-I DINO-I CLIP-T CLIP* DINO*
-GILL 0.55 0.49 0.43 0.23 0.21
-StackedDiffusion 0.76 0.46 0.47 0.36 0.22
-Image Sequence Decoding
-Greedy 0.81 0.54 0.48 0.39 0.26
-Nucleus 0.77 0.53 0.48 0.37 0.25
-BeamDiffusion 0.83 0.62 0.48 0.40 0.30
-5.3 Best Beam Search Configuration
-The hyperparameters beam width ($w$ in Eq.
-5) and steps back ($m$ in Eq.
-2), play a crucial role in the performance of BeamDiffusion.
-In this section, we report an exaustive evaluation on how different beam widths (2, 3, 4, and 6) and steps back (2, 3, and 4) improve the latent selection and the overall quality of the generation.
-A more detailed discussion on the impact of the latents selection is provided in Appendix C.
-We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix B.1 for details.
-To determine the best configuration, annotations were used to discard non-optimal configurations.
-The results in Table 4 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%.
-Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration.
-The configurations with beam width 6 performed the worst in all the settings tested.
-To assess the reliability of these findings, we used Gemini for the same evaluation on six tasks, achieving an 85% Fleiss’ Kappa agreement [6], indicating near-perfect alignment.
-The results from the extended Gemini evaluation (Table 4) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained the best, now achieving a 48.57% selection rate, a significant improvement over the 33.33% from the initial six-task evaluation.
-Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative.
-However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
-Table 3: Influence of different hyperparameter values according to human annotations.
-Steps Back Beam Width
-2 3 4 6
-2 8.33 8.33 25.00 8.33
-3 16.67 - - -
-4 16.67 8.33 8.33 -
-Table 4: Influence of different hyperparameter values according to Gemini annotations.
-Steps Back Beam Width
-2 3 4 6
-2 5.71 2.86 48.57 5.71
-3 28.57 - 2.86 -
-4 - - 2.86 -
+Table 3 compares all methods using metrics that measure sequence quality, including CLIP-I, DINO-I, CLIP-T, CLIP*, DINO*, and Goal Faithfulness [29], Step Faithfulness [29], and Cross-Image Consistency [29].
+BeamDiffusion achieves the highest combined scores (CLIP*: 40.00%, DINO*: 30.00%), reflecting strong performance in both visual and semantic consistency.
+This is further confirmed by faithfulness and consistency metrics, where BeamDiffusion leads in Goal Faithfulness (0.82), Step Faithfulness (0.92), and Cross-Image Consistency (0.61).
+StoryDiffusion performs well on CLIP-T (35.00%) and achieves high scores on the alignment metrics Goal Faithfulness: 0.78, Step Faithfulness: 0.87, Cross-Image Consistency: 0.61, showing strong semantic alignment and visual coherence despite lower combined metrics.
+1P1S achieves the highest DINO-I (30.00%) but lower CLIP-T (10.00%) and moderate alignment scores, indicating strong visual consistency but weaker semantic alignment.
+StackedDiffusion maintains balanced performance across CLIP-I, DINO-I, and CLIP* (around 25.00%) with moderate alignment scores, while GILL consistently underperforms across both automatic and alignment metrics.
+Overall, the relationships between CLIP-I/DINO-I and Cross-Image Consistency, and between CLIP-T and Goal/Step Faithfulness, show that the combined metrics CLIP* and DINO* reliably reflect both semantic and visual consistency.
+Together, these results demonstrate that BeamDiffusion is the most effective method for generating coherent and semantically faithful sequences of images.
 5.4 Non-Linear Sequence Denoising
 To analyze how BeamDiffusion explores the latent space, we measured the number of times that a sequence step provides latents for the generation of subsequent steps of the sequence.
-In this experiment we consider sequences of four steps, (Figure 2).
-Appendix C.2 provides an analysis at the level of the denoising iterations.
-Latents from step 1 were chosen 69.9% of
-Figure 2: Non-linear steps selection, according to selection opportunities.
-the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence.
-Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%.
+In this experiment we consider sequences of four steps.
+Latents from step 1 were chosen 69.9% of the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence.
+Selection rates decrease for later steps, with Step 3 at 46.8%, Step 2 at 33.3%, and the random seed at 15.6%.
 Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate.
@@
 This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate sequence steps.
+Appendix C.2 provides mode details.
+5.5 Effect of Latent History
+Figure 3: Visual comparison of sequences generated with cumulative latent access.
+The first row shows a sequence generated using latents only from step 1, the second row shows a sequence generated using latents from steps 1 through 5, and the bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
+To better understand the role of latent reuse, we compared three settings: using latents only from step 1, using latents from all steps (1–5), and using BeamDiffusion (Figure 3).
+Sequences generated from step 1 alone lacked detail and consistency, while access to all steps improved quality but came with higher computational cost.
+In contrast, BeamDiffusion, which relies only on the two most recent steps, achieved results comparable to or better than full access while being substantially more efficient.
+This efficiency arises because each new latent already encodes information from previous steps (Figure 2(b)), making full-history conditioning unnecessary.
+A more detailed analysis of intermediate cases, is provided in Appendix C.3.
 (a) Recipes domain task.
-(b) DIY task.
-(c) VIST story.
-Figure 3: Example tasks from three different domains: Recipes ("Jamaican Callaloo With Shrimp"), DIY:("How to Make a Door"), and VIST ("July 4th Barbecue").
+(b) VIST story.
+Figure 4: Example tasks from two different domains: Recipes ("Crema Catalana with salted caramel sauce") and VIST ("Fireworks - 4th July in Washington, DC").
 5.6 Qualitative Analysis
-To highlight how different decoding methods and models impact the semantic and visual consistency of generated outputs, we present several examples in Figures 3(a)3(c) and Appendix F.2.
-We compare three decoding methods (Greedy, Nucleus, and Beam) with two models (GILL and StackedDiffusion).
-In Figure 3(a), BeamDiffusion consistently maintains the same pan across all steps, a feature not always seen in Greedy or Nucleus methods.
-While GILL struggles to follow the step text and results in more inconsistencies, Beam performs better in preserving visual coherence compared to Greedy and Nucleus, though it still faces some challenges.
-For DIY tasks,as shown in Figure 3(b), BeamDiffusion can preserve basic elements such as the color of the plywood and a basic adherence to the text.
-However, overall, all methods face challenges in capturing precise details.
-In the VIST dataset, Figure 3(c), Greedy and Nucleus alternate between black-and-white and color images, while BeamDiffusion ensures color consistency throughout the sequence.
-Furthermore, background elements, such as bushes, are maintained in the first three steps, and the individuals in the first and third steps closely resemble each other.
-Overall, BeamDiffusion excels in maintaining both semantic and visual consistency, even in complex out-of-domain tasks, In contrast, methods such as StackedDiffusion and GILL encounter challenges in keeping a balance between text alignment and consistency across images.
+To highlight the impact of our approach on semantic and visual consistency, we present two examples, one from the Recipes domain (Figure 4(a)) and one from VIST (Figure 4(b)).
+In the recipe example, BeamDiffusion avoids major semantic errors and maintains consistent object shapes across steps; for instance, while StoryDiffusion and 1P1S show inconsistencies in the shape of the tins between steps 3 and 4, BeamDiffusion preserves them accurately.
+In the VIST example, BeamDiffusion maintains both semantic alignment and visual coherence throughout the sequence, accurately depicting two friends and preserving consistent colors and lighting across all steps.
+In contrast, baselines such as 1P1S introduce a temporal inconsistency, shifting from sunset to night and back to sunset, while StoryDiffusion misinterprets the step text by showing three individuals instead of two friends.
 6 Limitations
@@
 The theoretical complexity of BeamDiffusion is linear with the number of beams that we configure the model to explore.
-The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process.
+The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process, discussed in Appendix G.
 However, this provides a solid ground to explore exciting new ideas where stronger constraints can be enforced in the explored beams.
@@
 In this paper, we proposed a novel method to explore the latent diffusion space with a beam search strategy, to generate image sequences.
-BeamDiffusion is a theoretically sound framework, drawing on the strengths of text-based decoding techniques [40], with clear experimental gains over alternative methods, as we demonstrated in generating image sequences with enhanced coherence and consistency.
+BeamDiffusion is a theoretically sound framework, drawing on the strengths decoding techniques [44], with clear experimental gains over alternative methods, as we demonstrated in generating image sequences with enhanced coherence and consistency.
 Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences.
@@
 Solving this challenge, will improve the correlation between all elements in a sequence of scenes.
-Appendix A Implementation Details
-Base Model.
-Our method is implemented on top of Stable Diffusion v2.1 [35], a latent diffusion model for high-quality text-to-image synthesis.
-We build upon the open-source implementation provided by Stability AI, extending it to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text.
-Latent Trajectory Reuse.
-In our approach, maintaining coherence across sequential generations requires access to the intermediate latent representations produced during the generative process.
-Fortunately, the model pipeline inherently provides the full sequence of latent representations at each generation step.
-This enables us to easily extract and reuse specific latents as seed for new generations.
-Classifier Training.
-To train the contextual classifier, we employ a sequence-aware training approach that allows the model to capture dependencies between steps in the generation process.
-The classifier $\varphi$ takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and the respective step text ($B_{l-1}^{(b)}$).
-Then the classifier scores $I_{j}$ based on how well it fits within the evolving trajectory.
-During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$.
-The model is trained to minimize the cross-entropy loss:
-$\operatorname*{argmin}\sum_{d=1}^{D}\sum_{l=1}^{L}l_{d,l}\cdot\log(\varphi(I_{l},s_{l},B^{(b)}_{l-1})),$ (6)
-This contrastive training encourages the model to assign high scores to contextually appropriate continuations while penalizing incoherent or semantically unrelated ones.
-As a result, the classifier learns to distinguish subtle differences in step-to-step coherence, both visually and textually.
-Hardware.
-All experiments were ran on NVIDIA A100 GPUs with 40GB of memory.
-Appendix B Experiments Details
-B.1 Metrics and Annotations
+Appendix A Contextual Scene Descriptions
+In many cases, step descriptions alone may lack context as precise visual prompts.
+These step descriptions are often interdependent, relying on prior steps for contextual understanding, or may miss visual details, e.g.
+"Put the ice cream in freezer for 1 hour", or actions that involve multiple tasks, such as "Slice the tomatoes, grate the cheese, and stir the sauce".
+To address this challenge, we leverage Gemini [38], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context.
+Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
+$c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (6)
+Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps.
+Appendix B Best Beam Search Configuration
+The hyperparameters beam width ($w$ in Eq.
+5) and steps back ($m$ in Eq.
+2), play a crucial role in the performance of BeamDiffusion.
+In this section, we report an exaustive evaluation on how different beam widths (2, 3, 4, and 6) and steps back (2, 3, and 4) improve the latent selection and the overall quality of the generation.
+A more detailed discussion on the impact of the latents selection is provided in Appendix C.
+We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix E for details.
+To determine the best configuration, annotations were used to discard non-optimal configurations.
+The results in Table 5 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%.
+Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration.
+The configurations with beam width 6 performed the worst in all the settings tested.
+To verify the reliability of our findings, the LLM-as-a-judge was applied to six tasks, yielding an 85% Fleiss’ Kappa agreement [8], reflecting near-perfect alignment.
+The extended LLM-as-a-judge evaluation (Table 5) confirmed that the optimal configuration, 2 steps back with a beam width of 4, remained best, reaching a 48.57% selection rate compared to 33.33% in the initial six-task evaluation.
+Further, the configuration with 3 steps back and a beam width of 2 still performed well, yielding a 28.57% selection rate, confirming its robustness as an alternative.
+However, larger beam widths, such as beam width 6, continued to underperform, reinforcing that, in this task, a higher beam width does not always improve performance.
+Table 4: Influence of different hyperparameter values according to human annotations.
+Steps
+Back
+Beam Width
+2 3 4 6
+2 8.33 8.33 25.00 8.33
+3 16.67 - - -
+4 16.67 8.33 8.33 -
+Table 5: Influence of different hyperparameter values according to LLM-as-a-judge annotations.
+Steps
+Back
+Beam Width
+2 3 4 6
+2 5.71 2.86 48.57 5.71
+3 28.57 - 2.86 -
+4 - - 2.86 -
+Appendix C Impact of Latent Selection on Beam Search Performance
+Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search.
+By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence.
+This aligns with recent work on interpreting diffusion models.
+For example, Pont-Tuset et al.
+[34] ground textual descriptions in image regions, while Tang et al.
+[50] and Dewan et al.
+[7] explore how textual prompts influence generation through attention analysis and partial information decomposition.
+In contrast, our method contributes to interpretability by examining how latent space traversal, guided by beam search, affects the generation of coherent, semantically aligned image sequences.
+To systematically investigate this, we evaluate several sequences, each consisting of four steps, allowing us to assess the effects of latent selection across a diverse set of sequences.
+Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
+C.1 Impact of Latent Groups on Performance
+Following the initial analysis of beam search configurations (Appendix B), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences.
+While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance.
+In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and LLM-as-a-judge annotations.
+For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%.
+All other configurations resulted in 0% selection.
+In contrast, none of the configurations showed any non-zero selection rate under LLM-as-a-judge annotations, reinforcing the lack of preference for these latents.
+Figure 5: Impact of using late-index latents configurations on sequences generation.
+Similarly, for LLM-as-a-judge annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates.
+When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%.
+Figure 5 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [37, 3].
+C.2 Latent Selection on Image Sequence Generation
+Figure 6: The average number of times that the model selected a given latent to generate the next image.
+Following our evaluation of the impact of latent groups on performance, we now examine the role of individual latents within the higher-performing group (Latents 0-3).
+As shown in Figure 6, Latent 3 is selected most frequently (44.4%), followed by Latent 0 (20%), while Latents 1 and 2 are used less often (11.1% and 8.9%, respectively).
+This uneven distribution suggests that Latents 0 and 3 contribute disproportionately to the generation of sequences with better text alignment.
+One possible explanation is that these latents encode more semantically relevant or visually discriminative features that align well with the text prompts.
+Their frequent selection by beam search indicates that these latents are particularly effective at resolving ambiguities or reinforcing important visual cues during sequence generation.
+The lower usage of Latents 1 and 2 may reflect their tendency to encode less salient or redundant information, which contributes less to text-conditioned coherence.
+Interestingly, the Rand seed, which injects stochasticity into the generation process, accounts for 15.6% of selections.
+This suggests that while beam search primarily exploits strong latents like 0 and 3, it also values diversity.
+The inclusion of randomized elements allows the search to escape local optima and sample from alternative yet plausible paths, enhancing the robustness of the final sequence.
+C.3 Effect of Cumulative Latent Access
+To investigate how access to latents from previous steps influences the quality of generation, we conducted an ablation study in which we constrain the set of available latents throughout the generation of the sequence.
+Specifically, we evaluate performance when the model is limited to latents from only the first step, then from the first two steps, and so on, up to including latents from all previous steps (i.e., full cumulative access).
+Figure 7: Visual comparison of sequences generated with cumulative latent access.
+Each row shows a sequence generated using latents from only the first step, the first two steps, and so on, up to all previous steps.
+The bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
+As shown in Figure 7, including latents from only the first step makes step 4 appear very closed up.
+Limiting the model to latents from the first two steps makes it more difficult to follow the step text while maintaining coherence, which is particularly evident in step 5.
+Gradually expanding the latent set to include more steps improves sequence quality, especially when moving from two to three steps.
+However, including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
+In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case.
+The reason is that sequence generation naturally forms a latent chain: to generate a new image, the model conditions on a latent from the previous step, which has already absorbed information from earlier generations.
+The resulting latent is then updated with the new step, effectively carrying forward both old and new information (see Figure 2(b) for an illustration of this chain effect).
+In this way, the most recent latents act as a compressed summary of the entire history, making access to all earlier latents redundant.
+By always using only the last two latents, BeamDiffusion maintains a global summary of the sequence while preserving efficiency.
+Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary.
+Strategies such as BeamDiffusion, which leverage a limited yet informative temporal window, can produce high-quality sequences efficiently.
+These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents.
+C.4 Automatic and Qualitative Evaluation of Decoding Strategies
+Table 6: Comparison of different decoding methods across embedding-based sequence quality metrics.
+CLIP-I, DINO-I, CLIP-T, CLIP*, and DINO* are reported as win percentages, while Goal Faithfulness, Step Faithfulness, and Cross-Image Consistency are reported as absolute scores.
+Method CLIP-I DINO-I CLIP-T CLIP* DINO*
+[…]
+[…]
+[…]
+Greedy/CoSeDlen=1 30.00 30.00 35.00 25.00 30.00 0.76 0.90 0.62
+Nucleus Sampling 25.00 40.00 40.00 35.00 35.00 0.72 0.80 0.60
+BeamDiffusion 45.00 30.00 25.00 40.00 35.00 0.82 0.92 0.61
+Figure 8: Illustrative comparison of decoding approaches. BeamDiffusion prevents cascading inconsistencies across steps while preserving semantic and visual consistency.
+Automatic evaluation.
+Table 6 reports automatic comparisons across similarity metrics and faithfulness/consistency measures.
+The combined scores CLIP* and DINO* offer a more holistic signal than their individual components.
+Here, BeamDiffusion achieves the best CLIP* score (40.00) and matches the highest DINO* (30.00), indicating strong performance across both text–image and image–image similarity.
+By contrast, Greedy/CoSeD lags significantly on CLIP* (25.00) and only reaches 30.00 on DINO*, while Nucleus Sampling improves slightly on CLIP* (35.00) and DINO* (35.00) but at the expense of weaker instruction alignment and coherence.
+BeamDiffusion also obtains the strongest scores on goal faithfulness (0.82) and step faithfulness (0.92), showing that its generations remain closely aligned with both the overall task and individual step instructions.
+Greedy/CoSeD achieves reasonably high step faithfulness (0.90) and the best cross-image consistency (0.62), but falls behind in goal alignment.
+Nucleus Sampling, in contrast, is weaker across all three measures (0.72, 0.80, 0.60).
+Taken together, the similarity metrics and the faithfulness/consistency measures highlight that BeamDiffusion avoids the sharp trade-offs of the other decoding methods, achieving a favorable balance between multimodal similarity, instruction following, and visual coherence.
+Qualitative evaluation.
+Figures 8 and 9 illustrate characteristic error modes of baseline strategies.
+Nucleus Sampling often suffers from blur propagation, where low-frequency artifacts introduced in one step persist across subsequent generations, degrading the sequence.
+Greedy decoding avoids blur but instead produces visual inconsistencies, with objects abruptly shifting style or geometry across steps (e.g., the changing appearance of the pan between step 1 and step 3, in Figure 8).
+In contrast, BeamDiffusion avoids both pitfalls.
+By dynamically selecting informative latent trajectories, it prevents error accumulation while preserving smooth semantic and visual transitions.
+The generated sequences remain sharper, semantically faithful, and more coherent over time.
+This demonstrates that decoding is not merely a technical detail but a decisive factor in determining the fidelity and consistency of long-horizon generations.
+Figure 9: Illustrative comparison of decoding approaches. BeamDiffusion prevents cascading inconsistencies across steps while preserving semantic and visual consistency.
+Summary.
+Overall, both automatic and qualitative results show that BeamDiffusion delivers a principled improvement over standard strategies.
+By explicitly managing latent trajectories, it achieves robustness against error propagation and inconsistency while preserving semantic faithfulness and cross-step coherence.
+Appendix D Minimizing Decoding Risk
+The generation process using diffusion models begins with a seed that determines the initial conditions of the model.
+The choice of this seed significantly affects the outcome, as it influences the path the model takes during generation.
+Using multiple random seeds in the first step ensures a diverse exploration of possible outputs, avoiding premature narrowing of options into a single path based on that seed.
+This limits the diversity of the results and increases the chance of missing better options.
+Greedy Decoding and Nucleus Sampling.
+In both greedy decoding and nucleus sampling, we use the same approach in the first step.
+Multiple random seeds are employed to generate a set of candidate images.
+Each candidate is then evaluated using the CLIP score mechanism, which compares the generated images with the step text.
+The image that best aligns with the text, according to the CLIP score, is selected to continue the process.
+This method ensures that both greedy decoding and nucleus sampling benefit from multiple diverse starting points, increasing the quality and variety of the generated outputs.
+BeamDiffusion.
+In BeamDiffusion, relying on a single random seed in the first step effectively makes the process a greedy selection, as it locks the generation into a specific path.
+This limits diversity and reduces the chances of discovering higher-quality outputs.
+By using multiple random seeds, each seed generates a different decoding trajectory, helping to broaden the search space and increasing the likelihood of finding better solutions.
+Using multiple random seeds in the initial step enhances the exploration of the latent space, ensuring better and more diverse results across different decoding strategies.
+Random Seeds for mid-sequence Generation.
+The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$ may have one scene in the middle of the sequence that should be independent.
+To address this, for steps $s_{j>1}$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
+Figure 10: Illustration of the limitation of CLIP-I: despite a high CLIP-I score, the result is not semantically correct or visually faithful.
+The example compares several automatic metrics (CLIP-T, CLIP-I, DINO-I, CLIP*, DINO*), highlighting discrepancies and failure cases in sequence evaluation.
+Appendix E Metrics and Annotations
 Automatic Metrics.
@@
 [22] address this by clipping CLIP-T values below 0.1, effectively filtering out weak text-image matches.
-Figure 4: Illustration of the limitation of CLIP-I: despite a high CLIP-I score, the result is not semantically correct or visually faithful.
-The example compares several automatic metrics (CLIP-T, CLIP-I, DINO-I, CLIP*, DINO*), highlighting discrepancies and failure cases in sequence evaluation.
 To balance these factors, we set CLIP-I scores above 0.9 and DINO-I scores above 0.85 to zero, penalizing visual redundancy.
@@
 This example reinforces the importance of our clipping strategy: without penalizing excessively high CLIP-I values, the generation process may favor near-duplicates over meaningful visual progression, ultimately undermining semantic fidelity to the prompt.
-Table 5: Comparison of CLIP-I, DINO-I and CLIP-T across different methods with the standard deviation
-Method CLIP-I DINO-I CLIP-T
-GILL 0.55 $\pm$ 0.30 0.49 $\pm$ 0.26 0.43 $\pm$ 0.08
-StackedDiffusion 0.76 $\pm$ 0.12 0.46 $\pm$ 0.14 0.47 $\pm$ 0.09
-Image Sequence Decoding
-Greedy 0.81 $\pm$ 0.13 0.54 $\pm$ 0.14 0.48 $\pm$ 0.07
-Nucleus 0.77 $\pm$ 0.15 0.53 $\pm$ 0.15 0.48 $\pm$ 0.08
-BeamDiffusion 0.83 $\pm$ 0.14 0.62 $\pm$ 0.18 0.48 $\pm$ 0.09
-To complement average scores, we analyze the standard deviation ($\sigma$) of these metrics (Table 5).
-A low $\sigma$ reflects stable, consistent performance between diverse sequences, while a higher $\sigma$ implies variability in quality or adaptation to task complexity.
-GILL exhibits the highest $\sigma$ in CLIP-I and DINO-I, indicating its semantic similarity varies significantly between tasks, likely due to inconsistent adherence to step instructions.
-StackedDiffusion’s low $\sigma$ indicates its consistent performance across different tasks.
-BeamDiffusion strikes a balance with moderate $\sigma$ values, indicating a balance between consistency and variation across tasks.
-It is worth noting that CLIP-T generally shows lower standard deviation across models, since this metric is explicitly minimized during the diffusion model training.
-As a result, CLIP-T scores tend to be more stable regardless of the task-specific variation.
-Together, the mean and standard deviation metrics suggest BeamDiffusion offers the best combination of high semantic fidelity, text alignment, and stable adaptation across different sequences, making it the most robust method for sequential image generation.
+Even after applying this clipping, we observe that automatic metrics remain imperfectly aligned with human and LLM-as-a-judge evaluations, underscoring the ongoing challenge of developing reliable metrics for evaluating sequences of images.
 Human annotations.
@@
 Figure 12: Example of a human annotation to choose the best method.
-Gemini evaluation.
-Gemini’s selection process for annotating the best beam search configuration and the best generation method closely mirrored that of human annotators, with one key adjustment: Gemini tends to default to "both good" or "both bad" when such options are available.
-To ensure more decisive and accurate results, we removed these options from the prompt.
+LLM-as-a-judge Evaluation.
+The LLM-as-a-judge selection process for annotating the best beam search configuration and generation method closely mirrored that of human annotators, with one key adjustment: the model tended to default to "both good" or "both bad" when such options were available.
+To encourage more decisive outcomes, we removed these options from the prompt.
 We also refined our comparison methodology.
-Initially, presenting all five sequences at once led to biased evaluations, as the model disproportionately favored the last two sequences.
-To mitigate this, we adopted a pairwise evaluation strategy, comparing two sequences at a time.
-This approach resulted in more balanced and reliable assessments across all comparisons.
-Figure 7 shows the prompt used in Gemini to guide the evaluation of image sequences based on visual and semantic consistency.
-Gemini Prompt: "Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames.
+Presenting all five sequences simultaneously initially led to biased evaluations, as the model disproportionately favored the last two.
+To address this, we adopted a pairwise evaluation strategy, comparing two sequences at a time.
+This adjustment yielded more balanced and reliable assessments across all comparisons.
+Figure 13 presents the LLM-as-a-judge prompt for evaluating image sequences on visual and semantic consistency.
+LLM-as-a-judge Prompt:
+Evaluate two image sequences using the following criteria:
+Visual Consistency: Objects, ingredients, backgrounds, colors, and styles should remain consistent across frames.
 Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements.
 Semantic Consistency: Images must logically follow the described steps, clearly and accurately.
-Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic).
+Steps of the process: {steps} After evaluating both sequences, select the one with the strongest overall consistency (visual and semantic).
 Pay close attention to even minimal differences, as the sequences are very similar.
 If both sequences are equally good or bad, choose the one that is slightly better.
-Choose exactly one of the following options: - "First sequence" - "Second sequence" "
-Figure 7: Gemini prompt used to select the best beam search configuration and best method.
-B.2 Datasets
-Recipes and DIYs.
-Each manual task in both the Recipes and DIYs datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated.
+Choose exactly one of the following options:
+  - "First sequence"
+  - "Second sequence"
+Figure 13: LLM-as-a-judge prompt used to select the best beam search configuration and best method.
+Appendix F Datasets
+Recipes.
+Each manual task in the Recipes datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated.
 The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps.
 Most tasks feature an image for each step.
-The DIYs dataset includes a variety of tasks, from simpler projects to more complex ones, with detailed steps.
 VIST.
@@
 By using CLIP, we were able to select the annotation that best represented the image, improving the quality of our generated prompts and enhancing the overall alignment between text and visuals.
-Appendix C Impact of Latent Selection on Beam Search Performance
-Latent selection plays a crucial role in determining the quality and coherence of image sequences generated through beam search.
-By analyzing how different latent groups contribute to generation performance, we can better understand their impact on semantic alignment and visual coherence.
-This aligns with recent work on interpreting diffusion models.
-For example, Pont-Tuset et al.
-[30] ground textual descriptions in image regions, while Tang et al.
-[46] and Dewan et al.
-[5] explore how textual prompts influence generation through attention analysis and partial information decomposition.
-In contrast, our method contributes to interpretability by examining how latent space traversal, guided by beam search, affects the generation of coherent, semantically aligned image sequences.
-To systematically investigate this, we evaluate several sequences, each consisting of four steps, allowing us to assess the effects of latent selection across a diverse set of sequences.
-Using both automatic evaluation and human annotations, we analyze different latent configurations to determine which groups contribute most to generating coherent and semantically meaningful sequences.
-C.1 Impact of Latent Groups on Performance
-Annotations.
-Following the initial analysis of beam search configurations (Sec.
-5.3), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences.
-While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance.
-In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and Gemini annotations.
-For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%.
-All other configurations resulted in 0% selection.
-In contrast, none of the configurations showed any non-zero selection rate under Gemini annotations, reinforcing the lack of preference for these latents.
-Figure 8: Impact of using 1, 3, 5, and 7 latent configurations on image generation.
-Similarly, for Gemini annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates.
-When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%.
-Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [33, 2].
-Automatic Metrics.
-Given the observed negative impact of higher latents on performance in human annotations, we investigated how latents influence text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [46].
-DAAM interprets text-to-image diffusion models by analyzing cross-attention maps during the denoising process, generating heat maps that reveal how strongly each prompt token affects different regions of the output image.
-By applying DAAM, we examined whether higher latents distort or weaken the influence of key tokens during generation.
-The resulting heat maps allowed us to visually assess token-level contributions, helping us understand whether the model continues to attend to the most semantically important parts of the prompt under varying latent configurations..
-To quantify text alignment, we calculated the average heat map maximum intensity for all tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image.
-Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence.
-While this serves as an objective metric for evaluating text adherence, it remains a partial measure, as it focuses solely on text alignment without accounting for other important aspects of image quality, such as coherence, realism, or artistic fidelity.
-Figure 9 shows the average impact of different latent groups on text alignment.
-The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges.
-Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523).
-In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores.
-This suggests that earlier-stage latents play a crucial role in preserving text fidelity, whereas later-stage latents may dilute or distort the contribution of key tokens.
-Figure 9: The average influence of latent indices on text alignment across task data.
-In Figure 13, we analyze text alignment across multiple steps using heat maps corresponding to different latent groups, revealing how the connection between the input text and the generated image evolves throughout the process.
-Certain areas, such as the background and the bowl, show little to no heat map activation, indicating that these regions are less influenced by the text and more shaped by the model’s learned priors and overall scene composition.
-Since the model generates new images by building on latents from previous steps, these early decisions about text relevance persist throughout the process, influencing later refinements.
-Additionally, when examining generations using later latent groups, we find that this information remains more stable across all steps.
-This suggests that latents at these stages carry more structural and stylistic details, focusing on consistency rather than continuous adjustments based on the input text.
-While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts.
-For example, later latent groups (16-19) maintain stability in visual elements throughout the process, preserving previously established features.
-However, this comes at the expense of adapting to new or changing textual instructions.
-In contrast, early latent groups (0-3) demonstrate strong alignment between the text and the generated image, with consistently high heat map intensity across all steps.
-This highlights their role in grounding the image in the input text.
-For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
-C.2 Latent Selection on Image Sequence Generation
-Figure 10: The average number of times that the model selected a given latent to generate the next image.
-Following our evaluation of the impact of latent groups on performance, we now examine the role of individual latents within the higher-performing group (Latents 0-3).
-As shown in Figure 10, Latent 3 is selected most frequently (44.4%), followed by Latent 0 (20%), while Latents 1 and 2 are used less often (11.1% and 8.9%, respectively).
-This uneven distribution suggests that Latents 0 and 3 contribute disproportionately to the generation of sequences with better text alignment.
-One possible explanation is that these latents encode more semantically relevant or visually discriminative features that align well with the text prompts.
-Their frequent selection by beam search indicates that these latents are particularly effective at resolving ambiguities or reinforcing important visual cues during sequence generation.
-The lower usage of Latents 1 and 2 may reflect their tendency to encode less salient or redundant information, which contributes less to text-conditioned coherence.
-Interestingly, the Rand seed, which injects stochasticity into the generation process, accounts for 15.6% of selections.
-This suggests that while beam search primarily exploits strong latents like 0 and 3, it also values diversity.
-The inclusion of randomized elements allows the search to escape local optima and sample from alternative yet plausible paths, enhancing the robustness of the final sequence.
-Figure 11: Heat maps illustrating cross-attention in latent diffusion across different decoding methods.
-The visualizations show variations in how each approach preserves or redistributes attention to textual cues throughout the generation process.
-Figure 11 illustrates how different latent search methods affect text-image alignment.
-The heat maps show how attention patterns shift depending on the decoding strategy, revealing variations in how well each method preserves focus on key textual cues.
-Notably, in step 3, BeamDiffusion exhibits superior alignment, maintaining a stronger connection to the prompt and ensuring greater sequence coherence compared to other methods.
-While Beam Diffusion generally achieves better visual consistency and stronger text adherence in earlier steps, it slightly underperforms in step 5, failing to capture the "fireworks" mentioned in the prompt.
-This highlights how even methods with strong overall performance can exhibit localized weaknesses, underscoring the importance of analyzing alignment step by step.
-Overall, these findings highlight the significance of latent selection strategies in optimizing image generation.
-The results indicate that prioritizing the most significant latents could enhance the coherence of visual sequences and improve text alignment.
-These observations underline the influence of latent search strategies on the semantic accuracy and coherence of generated outputs.
-C.3 Effect of Cumulative Latent Access
-To investigate how access to latents from previous steps influences the quality of generation, we conducted an ablation study in which we constrain the set of available latents throughout the generation of the sequence.
-Specifically, we evaluate performance when the model is limited to latents from only the first step, then from the first two steps, and so on, up to including latents from all previous steps (i.e., full cumulative access).
-Figure 12: Visual comparison of sequences generated with cumulative latent access.
-Each row shows a sequence generated using latents from only the first step, the first two steps, and so on, up to all previous steps.
-The bottom row shows the result from BeamDiffusion, which uses latents from only the two preceding steps.
-As shown in Figure 12, limiting the model to latents from only the first step results in weak text-image alignment and incoherent visual transitions.
-Gradually expanding the latent set to include more steps improves sequence quality, particularly when moving from two to three steps.
-However, we observe that including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
-In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case.
-This indicates that a narrow temporal window is sufficient to maintain strong alignment and visual consistency.
-By avoiding the computational cost of considering all previous steps, BeamDiffusion finds a favorable trade-off between quality and efficiency.
-Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary.
-Strategies such as BeamDiffusion, which take advantage of a limited but informative temporal context, can produce high-quality sequences while significantly reducing computational overhead.
-These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents without the need for full sequence access.
-Appendix D Minimizing Decoding Risk
-The generation process using diffusion models begins with a seed that determines the initial conditions of the model.
-The choice of this seed significantly affects the outcome, as it influences the path the model takes during generation.
-Using multiple random seeds in the first step ensures a diverse exploration of possible outputs, avoiding premature narrowing of options into a single path based on that seed.
-This limits the diversity of the results and increases the chance of missing better options.
-Greedy Decoding and Nucleus Sampling.
-In both greedy decoding and nucleus sampling, we use the same approach in the first step.
-Multiple random seeds are employed to generate a set of candidate images.
-Each candidate is then evaluated using the CLIP score mechanism, which compares the generated images with the step text.
-The image that best aligns with the text, according to the CLIP score, is selected to continue the process.
-This method ensures that both greedy decoding and nucleus sampling benefit from multiple diverse starting points, increasing the quality and variety of the generated outputs.
-BeamDiffusion.
-In BeamDiffusion, relying on a single random seed in the first step effectively makes the process a greedy selection, as it locks the generation into a specific path.
-This limits diversity and reduces the chances of discovering higher-quality outputs.
-By using multiple random seeds, each seed generates a different decoding trajectory, helping to broaden the search space and increasing the likelihood of finding better solutions.
-Using multiple random seeds in the initial step enhances the exploration of the latent space, ensuring better and more diverse results across different decoding strategies.
-Random Seeds for mid-sequence Generation.
-The sequence of scenes $S=\{s_{1},s_{2},\dots,s_{L}\}$ may have one scene in the middle of the sequence that should be independent.
-To address this, for steps $s_{j>1}$, we use a combination of random seeds and past latent vectors to maintain a balance between diversity and coherence, which is particularly important for such scenarios.
-Appendix E Broader Impacts and Challenges
+Appendix G Challenges
 In the development and evaluation of the BeamDiffusion method, we encountered several challenges that required careful consideration and innovation.
@@
 In this section, we outline the key challenges we faced and the strategies we employed, while acknowledging that not all of them have been fully addressed.
-E.1 Impacts
-This work proposes a method to improve generative pretraining by incorporating subject-aware mutual information maximization, aiming to produce more contextually aligned and coherent sequences of images.
-The ability to generate visually consistent image sequences has positive applications in areas such as animation, visual storytelling, educational media, and accessibility tools (e.g., for generating illustrative content from text).
-However, as with many generative models, there are potential negative societal impacts.
-Enhanced coherence and subject fidelity could be misused to create misleading or deceptive visual content, such as more convincing visual disinformation, propaganda, or synthetic media (e.g., deepfakes).
-E.2 Contextual Scene Descriptions
-One of the challenges encountered concerns the generation of captions that accurately describe the steps involved in the sequence.
+Captions.
+One of the challenges encountered in this process is generating captions that accurately describe the steps involved in the sequence.
 Often, a single step may consist of multiple sub-steps, each contributing to the overall generation.
@@
 Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
-To address this challenge, we leverage Gemini [34], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context.
-Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
-$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (7)
-Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps.
-E.3 Domains.
+Domains.
 Identifying a domain that effectively highlights the advantages of the BeamDiffusion method, particularly the benefits of beam search, is still a challenge.
@@
 Finding a domain that not only aligns with the method’s capabilities but also provides a clear and compelling demonstration of its improvements remains an ongoing challenge.
-E.4 Complexity of the generation.
+Complexity of the generation.
 The BeamDiffusion model generates multiple candidate images at each step by exploring a range of latents.
@@
 While the initial three steps remain computationally demanding due to the expansive search space, pruning combined with this heuristic can significantly lower the computational cost by refining the candidate set.
-Appendix F Application Examples
+(a) Example 1: 50-step generations exhibit blur propagation, while 100 steps reduce blur and sharpen frames. 150 steps show no substantial gain over 100.
+(b) Example 2: 50 steps produce blurry, inconsistent frames, while 100 steps improve fidelity and coherence. 150 steps do not yield further noticeable improvements.
+Figure 14: Qualitative comparison of 50 vs. 100 denoising steps across two examples. Increasing steps to 150 does not produce substantial gains over 100, indicating diminishing returns.
+Appendix H Implementation Details
+Base Models.
+Our method is primarily implemented on top of FLUX.1-dev[2], a diffusion transformer–based model released by Black Forest Labs.
+We extend its open-source implementation to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text.
+Unless otherwise specified, we follow the default configuration of FLUX.1-dev with 100 denoising steps, chosen to balance generation quality and computational efficiency.
+While the default 50 steps generally produce globally coherent images, we observe that finer structures often remain underdeveloped, resulting in blur and reduced visual consistency across sequences.
+Increasing to 100 steps substantially sharpens details and improves sequence coherence, as illustrated in Figures 14(a) and 14(b).
+Further increasing to 150 steps provides no additional gains, confirming that 100 steps is a reasonable trade-off.
+To ensure that the choice of decoding method is not dependent on a specific backbone, we also implement BeamDiffusion on Stable Diffusion v2.1[39], a widely used U-Net–based latent diffusion model, using its standard parameters.
+This setup verifies that our conclusions regarding decoding strategies generalize across different generative architectures.
+Latent Trajectory Reuse.
+In our approach, maintaining coherence across sequential generations requires access to the intermediate latent representations produced during the generative process.
+Fortunately, the model pipeline inherently provides the full sequence of latent representations at each generation step.
+This enables us to easily extract and reuse specific latents as seed for new generations.
+Classifier Training.
+To train the contextual classifier, we employ a sequence-aware training approach that enables the model to capture dependencies between steps in the generation process.
+The classifier $\varphi$ is primarily trained on the Recipes dataset, which provides rich sequential image-text pairs of cooking steps, allowing it to learn meaningful context transitions in a structured procedural domain.
+Additionally, a small amount of data from the VIST dataset is included during training to introduce limited variability in narrative structure.
+The classifier takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and respective step text ($B_{l-1}^{(b)}$).
+It then scores $I_{l}$ based on how well it fits within the evolving trajectory.
+During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$. The model is trained to minimize the cross-entropy loss:
+$\operatorname*{argmin}\sum_{d=1}^{D}\sum_{l=1}^{L}l_{d,l}\cdot\log(\varphi(I_{l},s_{l},B^{(b)}_{l-1}))$, (7)
+This contrastive training encourages the model to assign high scores to contextually appropriate continuations while penalizing incoherent or semantically unrelated ones.
+As a result, the classifier learns to distinguish subtle differences in step-to-step coherence, both visually and textually.
+Hardware.
+All experiments were conducted on a NVIDIA A100 GPU with 40GB of memory.
+Appendix I Broader Impacts
+This work proposes a method to improve generative pretraining by incorporating subject-aware mutual information maximization, aiming to produce more contextually aligned and coherent sequences of images.
+The ability to generate visually consistent image sequences has positive applications in areas such as animation, visual storytelling, educational media, and accessibility tools (e.g., for generating illustrative content from text).
+However, as with many generative models, there are potential negative societal impacts.
+Enhanced coherence and subject fidelity could be misused to create misleading or deceptive visual content, such as more convincing visual disinformation, propaganda, or synthetic media (e.g., deepfakes).
+Appendix J Application Examples
 This section presents additional visual examples and qualitative results that complement the main paper.
@@
 In this appendix, we include expanded examples to further illustrate how our model addresses this issue.
-These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes, visual storytelling, and DIY tasks.
+These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes and visual storytelling.
 Together, these examples highlight the strengths and limitations of current approaches in achieving textual alignment and visual consistency in sequential image generation.
-Figure 13: Visualization of cross-attention in latent diffusion across different sets of latents.
-The heat maps reveal how attention patterns evolve at various sequence steps, highlighting the alignment between text prompts and distinct latent representations.
-F.1 Beam Search Tree
-(a) The beam denoising search tree.
-(b) Example of one complete decoded denoising process using contextual captions.
-Figure 14: The beam diffusion model works in the denoising latent space by denoising the initial seed to latents and them decoding different beam paths.
-This shows how our method evolves in the latent space and how it can explore the latent space.
-In Figure 14(a), we provide another example of how the latent search occurs within the beam denoising process.
-This illustrates the BeamDiffusion model at work, showing the denoising of the initial seed to latents and how different beam paths are decoded.
-The figure highlights how our method explores the latent space, iterating through various latents based on the beam search strategy.
-In Figure 14(b), we observe that the generation of the selected image at step 4 behaves similarly to an isolated diffusion process.
-However, it is important to note that this process is influenced by multiple prompts.
-Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4.
-This demonstrates how the latent search process incorporates context from various steps, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
-F.2 Sequence Examples
-To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 18-26.
-These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
-In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 18, 18, 22, and 22.
-They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps.
-This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt.
-In contrast, GILL and StackedDiffusion exhibit a different behavior.
-While both maintain high visual consistency, generating sequences in which the images are visually similar, they struggle to accurately follow the step-by-step textual instructions.
-GILL, in particular, tends to produce images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete.
-StackedDiffusion performs slightly better than GILL in this regard, capturing some of the step-wise changes, but still often fails to fully reflect the evolving textual input across the sequence.
-Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency.
-As shown in Figures 18, 18, 22, and 22, it generally produces the most coherent sequences compared to other methods.
-In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps.
-This combination of textual alignment and visual stability leads to more coherent and contextually accurate image sequences.
-While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches.
-However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 26, 26, and 26.
-The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence.
-While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
-Figure 15: Example of the Recipes domain task "Vegan Basil Pesto and Sun-Dried Tomato Pinwheels", with a sequence of 4 steps.
-Figure 16: Example of the Recipes domain task "Tangy lemon sheet cake", with a sequence of 4 steps.
-Figure 17: Example of the Recipes domain task "Zucchini, Pecorino & Mint Salad", with a sequence of 4 steps.
-Figure 18: Example of the Recipes domain task "Bone Broth", with a sequence of 3 steps.
-Figure 19: Example of the out-of-domain VIST story "Klamath Project Construction", with a sequence of 5 steps.
-Figure 20: Example of the out-of-domain VIST story "Yosemite Weekend", with a sequence of 5 steps.
-Figure 21: Example of the out-of-domain VIST story "Fireworks - 4th July in Washington, DC", with a sequence of 5 steps.
-Figure 22: Example of the out-of-domain VIST story "Independence Day Parade", with a sequence of 5 steps.
-Figure 23: Example of the out-of-domain DIY task "How to Clean Silk Rugs", with a sequence of 2 steps.
-Figure 24: Example of the out-of-domain DIY task "How to Polish Wood", with a sequence of 3 steps.
-Figure 25: Example of the out-of-domain DIY task "How to Preserve Basil", with a sequence of 6 steps.
-Figure 26: Example of the out-of-domain DIY task "How to Clean Natural Stone", with a sequence of 5 steps.
+J.1 Sequence Examples
+To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 16-18, highlighting how each method maintains sequence coherence in sequential image generation.
+For the Recipes domain ("Buttermilk chicken wings", Figure 16), most models incorrectly depict the chicken as already cooked in step 2, instead of starting raw and gradually appearing cooked.
+BeamDiffusion correctly captures this progression, faithfully reflecting the cooking process across all steps.
+Similarly, in the Recipes task "Sweet potato and orange soup" (Figure 16), GILL, StackedDiffusion, 1P1S, and StoryDiffusion misrepresent the ingredient being cut, while BeamDiffusion consistently depicts a person cutting a sweet potato and maintains stepwise visual consistency without hallucinations.
+For the VIST stories, BeamDiffusion demonstrates strong identity and style preservation.
+In "4th of July 2006" (Figure 18), it is the only method to maintain the same child between step 2 and step 4.
+In "Yosemite Weekend" (Figure 18), both BeamDiffusion variants maintain coherent tree styles across steps, whereas other models show a noticeable break between step 1 and step 2.
+Overall, while baseline methods may partially adhere to textual instructions or maintain visual similarity, BeamDiffusion consistently balances stepwise textual alignment and visual coherence, producing sequences that are both contextually accurate and visually stabl
+Figure 15: Example of the Recipes domain task "Buttermilk chicken wings", with a sequence of 4 steps.
+Figure 16: Example of the Recipes domain task "Sweet potato and orange soup", with a sequence of 4 steps.
+Figure 17: Example of the VIST story domain "4th of July 2006", with a sequence of 5 steps.
+Figure 18: Example of the VIST story domain "Yosemite Weekend", with a sequence of 5 steps.